### v2

Abstract
In this paper, we study the risk sharing problem among multiple agents using Lambda Value-at-Risk as their preference functional, under heterogeneous beliefs, where beliefs are represented by several probability measures. We obtain semi-explicit formulas for the inf-convolution of multiple Lambda Value-at-Risk measures under heterogeneous beliefs and the explicit forms of the corresponding optimal allocations. To show the impact of belief heterogeneity, we consider three cases: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs. For those cases, we find more explicit expressions for the inf-convolution, showing the influence of the relation of the beliefs on the inf-convolution. Moreover, we consider, in a two-agent setting, the inf-convolution of one Lambda Value-at-Risk and a general risk measure, including expected utility, distortion risk measures and Lambda Value-at-Risk as special cases, with differing beliefs. The expression of the inf-convolution and the form of the optimal allocation are obtained. In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high. Finally, we discuss risk sharing for an alternative definition of Lambda Value-at-Risk.
Key-words: Lambda Value-at-Risk; Value-at-Risk; Risk sharing; Inf-convolution; Distortion risk measure; Expected shortfall; $\mathrm{CoVaR}$; $\mathrm{CoES}$
1 Introduction
The Pareto-optimal risk sharing problem, with the preferences of agents represented by risk measures, has been studied extensively since the introduction of convex and coherent risk measures by Artzner et al. (1999), Föllmer and Schied (2002) and Frittelli and Rosazza Gianin (2005). For instance, convex risk measures were used in Barrieu and El Karoui (2005), Jouini et al. (2008) and Filipović and Svindland (2008) to investigate the risk sharing problem, showing the existence of the optimal allocation, which is comonotonic under the assumption of law-invariance. Recently, more focus has been placed on the risk sharing problem for non-convex risk measures, such as quantile-based risk measures including Value-at-Risk ($\mathrm{VaR}$) and Expected Shortfall ($\mathrm{ES}$) as special cases (e.g., Embrechts et al. (2018), Liu et al. (2022) and Weber (2018)) and distortion riskmetrics including inter-quantile-range as an example (e.g., Lauzier et al. (2023b)). In those papers, it is shown that the optimal allocation is pairwise countermonotonic, as defined in Lauzier et al. (2023a).
In the above papers, the agents are assumed to share the same beliefs on the distribution of the future risk. However, in the current regulatory frameworks (e.g., BCBS Standards (2016)), internal models are extensively used, leading to heterogeneous views of different agents regarding the same future risk. Moreover, the belief heterogeneity may also stem from the asymmetric information accessed by the agents. We refer to Embrechts (2017) for a discussion of application of internal models in banking and insurance and Xiong (2013) for a relevant discussion in finance. The model/belief heterogeneity in the risk sharing problem has been considered by several papers in the literature. For instance, under model heterogeneity, the risk sharing between two agents endowed with concave monetary utilities was studied in Acciaio and Svindland (2009), where the random variables were constrained to a finite $\sigma$-algebra and the existence of Pareto optimal allocation was discussed. Risk sharing for $\mathrm{VaR}$ or $\mathrm{ES}$ under heterogeneous beliefs was studied in Embrechts et al. (2020), who showed that optimal risk allocations are pairwise countermonotonic with model heterogeneity. Liu (2020) investigated the comonotonic risk sharing with heterogeneous beliefs for distortion risk measures and Liebrich (2024) studied the risk sharing for consistent risk measures with model heterogeneity, obtaining a sufficient condition for the existence of the optimal allocations under some assumptions on the relation of the beliefs. Model heterogeneity was also studied in the context of optimal insurance and reinsurance design; see Amarante et al. (2015), Boonen (2016), Chi (2019), Boonen and Ghossoub (2020), Asimit et al. (2021) and the references therein.
In this paper, we use Lambda Value-at-Risk ($\Lambda\mathrm{VaR}$) to represent the agents’ preferences. As an extension of $\mathrm{VaR}$, $\Lambda\mathrm{VaR}$ was introduced by Frittelli et al. (2014) by changing the fixed probability level to a probability/loss function $1-\Lambda$. The $\Lambda$ function can be chosen either increasing (relatively risk-averse) or decreasing (relatively risk-seeking), representing decision makers’ individual risk appetite, as shown in Frittelli et al. (2014). The choice of $\Lambda$ function based on data was studied in Hitaj et al. (2018). Compared with $\mathrm{VaR}$, one advantage of Lambda Value-at-Risk is its ability to distinguish the tail risk for decreasing $\Lambda$ functions, in the same spirit as the Loss $\mathrm{VaR}$ proposed in Bignozzi et al. (2020). For increasing $\Lambda$ functions, $\Lambda\mathrm{VaR}$ may incorporate some additional requirement such as risk manager’s judgement in the process of risk management; see Bellini and Peri (2022). Moreover, the recent literature shows that $\Lambda\mathrm{VaR}$ satisfies other desirable properties. Han et al. (2024) showed that $\Lambda\mathrm{VaR}$ with increasing $\Lambda$ function satisfies quasi-star-shapedness, which is a property weaker than quasi-convexity and penalizing a kind of risk concentration. In addition, $\Lambda\mathrm{VaR}$ with increasing $\Lambda$ functions also satisfies cash subadditivity, which is useful to measure the future financial loss with stochastic interest rates; see El Karoui and Ravanelli (2009) and Han et al. (2024). We refer to Bellini and Peri (2022) for the monotonicity, locality and other properties, and Burzoni et al. (2017) for robustness, elicitability and consistency. The risk sharing problem for $\Lambda\mathrm{VaR}$ under homogeneous beliefs was studied in Liu (2024) and Xia and Hu (2024), where the expressions for the inf-convolution and the forms of the optimal allocation were derived. The $\Lambda\mathrm{VaR}$ was also applied to optimal reinsurance design with model homogeneity in Balbás et al. (2023) and Boonen et al. (2024). Moreover, the application of $\Lambda\mathrm{VaR}$ to robust portfolio selection and sensitivity analysis can be found in Han and Liu (2024) and Ince et al. (2022).
In our study, we put our focus on the risk sharing problem with $\Lambda\mathrm{VaR}$ and belief heterogeneity, extending many results in the literature. In Section 3, we study the inf-convolution of multiple $\Lambda\mathrm{VaR}$ with heterogeneous beliefs, where the beliefs are represented by a set of probability measures. We obtain a semi-explicit formula for the inf-convolution and explicit forms for the optimal risk allocations, which show that agents take contingent losses on disjoint sets of the sample space.
The relation between the beliefs plays a crucial role in the risk sharing problem; see e.g., Chi (2019). To study the impact of belief heterogeneity, we consider three cases in Section 3: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs. For homogeneous beliefs, we show that the inf-convolution of $\Lambda\mathrm{VaR}$ is still a $\Lambda\mathrm{VaR}$ for general $\Lambda$ functions, extending the results in Liu (2024) and Xia and Hu (2024), where the monotonicity of the $\Lambda$ functions is typically required. In particular, our results include the case that some $\Lambda\mathrm{VaR}$ have increasing $\Lambda$ functions and others have decreasing $\Lambda$ functions, demonstrating that the agents may have relatively different risk appetites. The conditional beliefs reflect the asymmetric information obtained by the agents such that each agent has slightly different concern. Under this setup, we also obtain an explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$, showing that the inf-convolution of $\Lambda\mathrm{VaR}$ under conditional beliefs is also a $\Lambda\mathrm{VaR}$ under a new conditional belief. Conditional beliefs are closely related to the conditional risk measures proposed in the literature to measure systemic risk, such as $\mathrm{CoVaR}$ (Adrian and Brunnermeier (2016) and Girardi and Tolga Ergün (2013)) and $\mathrm{CoES}$ (Mainik and Schaanning (2014)). The risk sharing problem in this case can be interpreted as risk sharing with some Co-risk measures ($\mathrm{Co\Lambda\mathrm{VaR}}$, a new conditional risk measure proposed in Section 3). For the third case, we consider the risk sharing between two agents and suppose one probability measure is absolutely continuous with another one. Under this assumption, we find a more explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$. Note that we also give the results for the inf-convolution of $\mathrm{VaR}$ with conditional beliefs and absolutely continuous beliefs respectively, which are also new to the literature.
In Section 4, we study the inf-convolution of one $\Lambda\mathrm{VaR}$ and a general monotone risk measure with belief heterogeneity. The expression of the inf-convolution and the form of the optimal allocation are derived. Then we consider two cases: conditional beliefs and absolutely continuous beliefs. For conditional beliefs, we obtain explicit expressions for the inf-convolution of one $\Lambda\mathrm{VaR}$ and one distortion risk measure/expected utility/$\Lambda\mathrm{VaR}^{+}$. For absolutely continuous beliefs, we find the formula for the inf-convolution of one $\Lambda\mathrm{VaR}$ and one $\mathrm{ES}$/expected utility. In Section 5, we consider the inf-convolution of one $\Lambda\mathrm{VaR}^{+}$ (an alternative definition of Lambda Value-at-Risk) and a general monotone risk measure with belief heterogeneity. We obtain the expression of the inf-convolution, which is complicated, and the forms of the optimal allocation. Our results include the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs as a special case. The results here are not explicit and very different from the previous sections, as they involve the shape of the $\Lambda$ function and the best case of risk aggregation under dependence uncertainty. This means that $\Lambda\mathrm{VaR}^{+}$ is very different from $\Lambda\mathrm{VaR}$ for its application within the risk sharing problem.
Throughout the paper, the stated results provide conditions for trivial, i.e., infinite, inf-convolutions. We consistently show that these conditions arise when, broadly speaking, agents’ beliefs are strongly inconsistent and risk tolerances are high. Hence, our contribution illuminates how the entanglement of beliefs and preferences impacts key features of the risk sharing problem with $\Lambda\mathrm{VaR}$.
The notation and key definitions are given in Section 2 and all the proofs are delegated to Appendix A-D.
2 Notation and Definitions
For a given atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $L^{\infty}$ denote the collection of all bounded random variables. Moreover, let $\mathcal{X}$ be a set of random variables containing $L^{\infty}$. We suppose $\mathcal{X}$ has good enough properties to conduct our study, such as $\mathcal{X}=L^{p}$ for some $p\geqslant 0$, where $L^{p}$ represents the set of all random variables with finite $p$-th moments. Throughout this paper, we assume that the probability measures $\mathbb{Q},\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ live on $(\Omega,\mathcal{F})$ and the corresponding probability spaces are atomless. For any $X\in\mathcal{X}$, a positive value of $X$ represents a financial loss. The distribution function of $X$ under $\mathbb{Q}$ is denoted as $F_{X}^{\mathbb{Q}}$. For a mapping $\rho:\mathcal{X}\to\mathbb{R}$, we say $\rho$ is law-invariant under $\mathbb{Q}$ if for all $X,Y\in\mathcal{X}$,
$X\overset{\mathbb{Q}}{=}Y~{}~{}\Rightarrow~{}~{}\rho(X)=\rho(Y),$ (1)
where $\overset{\mathbb{Q}}{=}$ stands for equality in distribution under $\mathbb{Q}$; $\rho$ is monotone if $X\leqslant Y$ implies $\rho(X)\leqslant\rho(Y)$; and $\rho$ is cash-additive if $\rho(X+c)=\rho(X)+c$ for $X\in\mathcal{X}$ and $c\in\mathbb{R}$. We say $\rho$ is a monetary risk measure if it is monotone and cash-additive. For more details on risk measures, one can refer to Chapter 4 of Föllmer and Schied (2016). For simplicity, throughout the paper, we use the notation $\rho^{\mathbb{Q}}$ to represent that $\rho$ is law-invariant under $\mathbb{Q}$. For $X\in\mathcal{X}$, $F_{X}^{\mathbb{Q},-1}$ represents its left-quantile under $\mathbb{Q}$, which is defined by
$F_{X}^{\mathbb{Q},-1}(p)=\inf\{x:F_{X}^{\mathbb{Q}}(x)\geqslant p\},~{}p\in(0,1]$
with the convention that $\inf\emptyset=\infty$. For any $X\in\mathcal{X}$, we denote by $U_{X}^{\mathbb{Q}}$ a uniform random variable on $[0,1]$ under $\mathbb{Q}$ such that $X=F_{X}^{\mathbb{Q},-1}(U_{X}^{\mathbb{Q}})$ a.s. under $\mathbb{Q}$. The existence of such $U_{X}^{\mathbb{Q}}$ for any random variable $X$ is guaranteed by e.g., Lemma A.32 of Föllmer and Schied (2016).
Next, we define the inf-convolution. The random variable $X\in\mathcal{X}$ represents a risk to be shared between $n$ agents. Then, define the set of allocations of $X$ as
$\mathbb{A}_{n}(X)=\left\{(X_{1},\ldots,X_{n})\in\mathcal{X}^{n}:\sum_{i=1}^{n}X_{i}=X\right\}.$
In this paper, we suppose the agents may have heterogeneous beliefs, represented by some probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$. The inf-convolution of the risk measures $\rho_{1}^{\mathbb{Q}_{1}},\dots,\rho_{n}^{\mathbb{Q}_{n}}$ is the mapping $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}:\mathcal{X}\to[-\infty,\infty)$, defined as
$\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)=\inf\left\{\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i}):(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)\right\}.$
Note that $\rho_{i}$ is law-invariant under $\mathbb{Q}_{i}$, where $\mathbb{Q}_{i}$ represents the belief of the i-th agent.
An $n$-tuple $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is called an optimal allocation of $X$ for $(\rho_{1}^{\mathbb{Q}_{1}},\dots,\rho_{n}^{\mathbb{Q}_{n}})$ if $\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i})=\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$. The inf-convolution $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}(X)$ can be interpreted as the smallest possible aggregate capital for the total risk $X$ in financial system, if $\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ represents the capital charge for the i-th financial institution to hold the risky position $X_{i}$. More economic interpretations on the inf-convolution can be found in e.g., Delbaen (2012), Rüschendorf (2013) and Embrechts et al. (2018).
For the risk measures $(\rho_{1},\dots,\rho_{n})$ and a total risk $X$, an allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is Pareto-optimal if for any other allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$, $\rho_{i}(Y_{i})\leqslant\rho_{i}(X_{i})$ for all $i=1,\dots,n$ implies $\rho_{i}(Y_{i})=\rho_{i}(X_{i})$ for all $i=1,\dots,n$. For finite-valued monetary risk measures, it is shown in Embrechts et al. (2018) that an allocation is optimal if and only if it is Pareto-optimal. Note that $\Lambda\mathrm{VaR}$ (defined in (2)) does not satisfy cash-additivity; see Frittelli et al. (2014) and Bellini and Peri (2022). Hence it is not a monetary risk measure. However, one can still show that the optimal allocation of the inf-convolution of $\Lambda\mathrm{VaR}$ is Pareto-optimal.
Finally, we define the Lambda Value-at-Risk. For $\Lambda:\mathbb{R}\to[0,1]$ and a probability measure $\mathbb{Q}$, the Lambda Value-at-Risk are given by
$\Lambda\mathrm{VaR}^{\mathbb{Q}}(X)$ $=\inf\{x\in\mathbb{R}:F_{X}^{\mathbb{Q}}(x)\geqslant 1-\Lambda(x)\},$
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}(X)$ $=\sup\{x\in\mathbb{R}:F_{X}^{\mathbb{Q}}(x)<1-\Lambda(x)\},$ (2)
where $\inf\emptyset=\infty$ and $\sup\emptyset=-\infty$. Note that $\Lambda\mathrm{VaR}(X)=\Lambda\mathrm{VaR}^{+}(X)$ if $\Lambda$ is increasing; Otherwise, it may not be true; see Proposition 6 of Bellini and Peri (2022). We refer to Bellini and Peri (2022) for two other definitions of $\Lambda\mathrm{VaR}$. If $\Lambda$ is a constant, then $\Lambda\mathrm{VaR}^{\mathbb{Q}}$ boils down to $\mathrm{VaR}^{\mathbb{Q}}$, i.e., $\mathrm{VaR}^{\mathbb{Q}}$ at level $p\in[0,1)$ is given by
$\mathrm{VaR}_{p}^{\mathbb{Q}}(X)=\Lambda\mathrm{VaR}^{\mathbb{Q}}(X)=F_{X}^{\mathbb{Q},-1}(1-p),~{}~{}~{}~{}X\in\mathcal{X},$
for $\Lambda=p$. Moreover, by definition, $\mathrm{VaR}_{1}^{\mathbb{Q}}(X)=-\infty$. Although $\mathrm{VaR}$ has been criticized from different angles, it has been widely applied in practice for risk management due to its simplicity and possession of some nice properties; see McNeil et al. (2015) and the references therein for more detailed discussion on $\mathrm{VaR}$. Compared to $\mathrm{VaR}$, $\Lambda\mathrm{VaR}$ is more flexible in the choice of $\Lambda$ functions; it satisfies cash subadditivity and quasi-star-shapedness for increasing $\Lambda$; and $\Lambda\mathrm{VaR}^{+}$ is able to capture the tail risk for decreasing $\Lambda$ functions; see e.g., Frittelli et al. (2014), Hitaj et al. (2018) and Han et al. (2024).
We denote by $\mathcal{H}$ the collection of all right-continuous functions $\Lambda:\mathbb{R}\to[0,1]$. Hereafter, for any $\Lambda$, we denote $\lambda^{-}=\inf_{x\in\mathbb{R}}\Lambda(x)$ and $\lambda^{+}=\sup_{x\in\mathbb{R}}\Lambda(x)$. We say that a constant $\lambda$ is attainable for $\Lambda$ if there exists $x\in\mathbb{R}$ such that $\Lambda(x)=\lambda$.
The interplay among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ plays an important role in the inf-convolution, which is our main concern in this paper. This can be seen from the following proposition. We say a mapping $\rho:\mathcal{X}\to\mathbb{R}$ is constant-preserving if $\rho(c)=c$ for all $c\in\mathbb{R}$. Note that both $\Lambda\mathrm{VaR}^{\mathbb{Q}}$ and $\Lambda\mathrm{VaR}^{+,\mathbb{Q}}$ satisfy constant preservation.
Proposition 1.
Suppose $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ are mutually singular, and $\rho_{2}$ is constant-preserving. Then for $X\in\mathcal{X}$,
$\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)=-\infty.$
The conclusion in Proposition 1 implies that for any two monetary risk measures $\rho_{1}$ and $\rho_{2}$, $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)=-\infty$ if $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ are mutually singular. This means that if the beliefs of agents are too divergent, then each one of them can get unreasonably good deal by transferring risks between them. Then, the fact that the inf-convolution is $\infty$- ∞ reflects that agents can always make their position better and hence there is no optimal allocation. In contrast, if $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)>-\infty$ for many monetary risk measures $\rho_{1}$ and $\rho_{2}$; see e.g., Embrechts et al. (2018) for quantile-based risk measures and Filipović and Svindland (2008) for convex risk measures. Note that the conclusion in Proposition 1 can be easily extended to the case with $n$ agents for $n\geqslant 3$.
3 Inf-convolution of multiple $\Lambda\mathrm{VaR}$
In this section, we investigate the inf-convolution of $n$ Lambda Value-at-Risk with heterogeneous beliefs represented by $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$, respectively. Suppose $\mathcal{X}\supseteq L^{\infty}$. We find an expression for the inf-convolution and also the corresponding optimal allocation, covering the results of the inf-convolution of $\Lambda\mathrm{VaR}$ or $\mathrm{VaR}$ with homogeneous or heterogenous beliefs in Liu (2024), Embrechts et al. (2020) and Embrechts et al. (2018). We first consider the general beliefs and then discuss three different cases about the relation of the beliefs: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs.
3.1 General beliefs
In this subsection, we suppose $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are general probability measures. We first introduce the following notation. Let $\Pi_{n}(\Omega)=\{(A_{1},\dots,A_{n}):\cup_{i=1}^{n}A_{i}=\Omega,~{}A_{i}\cap A_{j}=\emptyset,~{}i\neq j\}$. For $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, let
$\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Note that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)\in[-\infty,\infty)$.
Theorem 1.
For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X).$
Moreover, if the minimizer of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$ exists denoted by $(y_{1}^{*},\dots,y_{n}^{*},A_{1}^{*},\dots,A_{n}^{*})$, then the optimal allocation is given by
$X_{i}=\left(X-\sum_{i=1}^{n}y_{i}^{*}\right)\mathds{1}_{A_{i}^{*}}+y_{i}^{*},~{}i=1,\dots,n.$ (3)
The allocation (3) can be understood as follows. The $i$-the agent is allocated a constant loss $y_{i}^{*}$ plus a contingent loss, which is the excess of the total risk over the aggregate cash allocation, restricted to the set $A_{i}^{*}$. The key feature of those sets is seen by the definition of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$: for each agent, their joint probability of (i) the set $A_{i}^{*}$ on which they suffer the contingent loss and (ii) the total loss exceeding the aggregate cash allocation, is restricted by their $\Lambda_{i}$ function.
Note that Theorem 1 is very general, covering Theorems 1, 2, 4 of Liu (2024) and Theorem 4 of Embrechts et al. (2020) as special cases. Compared with the results in Liu (2024), our results in Theorem 1 offer semi-explicit formulas for the inf-convolution. Later, we will specify the relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ to obtain more explicit formulas for the inf-convolution. Moreover, Theorem 1 does not explicitly show when the inf-convolution is $\infty$- ∞. The discussion on the finiteness of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$ is complicated as it heavily relies on the properties of Lambda functions and the relationship between $\mathbb{Q}_{i}$; see Liu (2024) and Proposition 1 in Section 2. This will be discussed in the next two propositions. Finally, note that if all $\Lambda_{i}$ are constants, then Theorem 1 boils down to Theorem 4 of Embrechts et al. (2020), where the inf-convolution of $\mathrm{VaR}$ has been studied with belief heterogeneity.
Next, let us discuss the finiteness of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. Let $x\wedge y=\min(x,y)$, $x\vee y=\max(x,y)$, $\bigwedge_{i=1}^{n}x_{i}=\min_{i=1}^{n}x_{i}$ and $\bigvee_{i=1}^{n}x_{i}=\max_{i=1}^{n}x_{i}$.
Proposition 2.
Let $\mathcal{X}=L^{\infty}$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
  1. (i)
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\leqslant 1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$;
  2. (ii)
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ for all $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
Note that if all $\Lambda_{i}$ are constants, then Proposition 2 offers a sufficient and necessary condition for the finiteness of the inf-convolution. In general, due to the complexity of the $\Lambda_{i}$ functions, those conditions are sufficient but not necessary.
As we can see from Proposition 2, cases (i) and (ii) cannot cover all scenarios. The discussion of other cases may require more information on the Lambda functions. We next consider all monotone Lambda functions in the same direction.
Proposition 3.
Let $\mathcal{X}=L^{\infty}$. Suppose all $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 are decreasing. Then we have the following conclusion.
  1. (i)
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$;
  2. (ii)
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ for all $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
Suppose all $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 are increasing. Then we have the following conclusion.
  1. (i)
    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)<1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$;
  2. (ii)
    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)>1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
In this more refined result, we see that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ plays a pivotal role when $\Lambda$ functions are decreasing. We see that, in order to guarantee a fininte inf-convolution for decreasing $\Lambda_{i}$, it is necessary for this quantity to exceed 1. For simplicity, assume that for all $i$, $\lambda_{i}^{+}$ is attained by $\Lambda_{i}$. Then, the condition in case (ii) of Proposition 3 for deceasing $\Lambda_{i}$ is violated when one can find a partition of $\Omega$ such that for all agents we have $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{+}$. This means that for each agent we can find a set $A_{i}$ on which the risk is suitably bounded by the largest acceptable probability to the agent of an adverse tail event. Motivated by this, we may consider $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ as a measure of belief homogeneity, which also reflects preferences. Consider two extreme cases. First, if sets $A_{1},\dots,A_{n}$ exist for which we have $\mathbb{Q}_{i}(A_{i})=0$ for all $i$, then we will have $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}=0.$ The ability to find such sets is of course a case of extremely diverging beliefs. On the other hand, consider the case of belief homogeneity, $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. Then, we will have $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ if and only if $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, which reflects the impact of preferences. Clearly, when the values of $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, indicating that agents accept lower total probabilities of adverse events, the inf-convolution is finite. Whereas, the higher risk tolerance of the tail events (i.e., $\sum_{i=1}^{n}\lambda_{i}^{+}>1$) enables the possibility of finding a trivially good deal for all agents, leading to an infinite inf-convolution.
When the functions are increasing, the picture is somewhat more complicated. Assuming attainability of all $\lambda_{i}^{-},\lambda_{i}^{+}$, the condition of part (i) is satisfied if we can find a partition for which, for some $i$, we have that $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{-}$ and $\mathbb{Q}_{j}(A_{j})<\lambda_{j}^{+},~{}j\neq i$. Hence, the conditions of Proposition 3 compared to those in Proposition 2, are weaker for part (i) and essentially weaker for part (ii), as the effect of the additional assumption of the $\Lambda$ functions.
3.2 Homogeneous beliefs
In this subsection, we focus on the case $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. We find more explicit expression for the inf-convolution. Let $\Lambda^{*}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right)$ for all $x\in\mathbb{R}$. We say $\Lambda^{*}$ is attainable if for each $x\in\mathbb{R}$, there exists $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ satisfying $\sum_{i=1}^{n}y_{i}=x$ such that $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})=\Lambda^{*}(x)$. For homogeneous beliefs, $\Lambda^{*}(x)$ represents the largest aggregate risk tolerance of the agents to control the tail event $\{X>x\}$ in the definition of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
Theorem 2.
For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, if $\Lambda^{*}$ is attainable, then we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X).$
Moreover, if additionally $x^{*}=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)>-\infty$ and $\Lambda^{*}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(X>x^{*},A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
Note that Theorem 2 shows that the optimal sets $(A_{1}^{*},\dots,A_{n}^{*})$ may not be unique. Under the assumption of Theorem 2, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$, optimal sets can be chosen as
$A_{i}^{*}$ =$=\left\{1-\sum_{j=1}^{i}\Lambda_{j}(y_{j}^{*})<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{i-1}\Lambda_{j}(y_{j}^{*})\right\},~{}i=1,\dots,n-1,$
$A_{n}^{*}$ =$=\left\{\left(1-\sum_{j=1}^{n}\Lambda_{j}(y_{j}^{*})\right)\wedge 0<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{n-1}\Lambda_{j}(y_{j}^{*})\right\},$
with the convention that $\sum_{j=1}^{0}a_{i}=0$. Moreover, the right-continuity of $\Lambda^{*}$ at $x^{*}$ can be removed if all $\Lambda_{i}$ are decreasing. Given the non-uniqueness of the allocation and the particular form of $A_{i}^{*}$, it is clear that agents’ preferences do not impact which specific part of the (tail of) $X$ they are exposed to at the optimum. What is however impacted is the probability of the contingent events, specifically, for $\Lambda^{*}(x^{*})<1$, we have $\mathbb{P}(A_{i}^{*})=\Lambda_{i}(y_{i}^{*}),~{}i=1,\dots,n-1$ and $\mathbb{P}(A_{n}^{*})=1-\sum_{j=1}^{n-1}\Lambda_{j}(y_{j}^{*}).$
Note that in Theorem 2, we do not require any monotonic properties of Lambda functions. The Lambda functions are very general with only one requirement on the attainability of $\Lambda^{*}$. Hence, it extends the results of Theorems 1, 2, 4 of Liu (2024), who investigates the inf-convolution of multiple $\Lambda\mathrm{VaR}$ with monotone Lambda functions for homogeneous beliefs. In particular, Theorem 2 covers the case that some $\Lambda\mathrm{VaR}$ have increasing $\Lambda$ functions and others have decreasing $\Lambda$ functions, demonstrating that the agents may have relatively different risk appetites.
We can also remove the assumption on the attainability of $\Lambda^{*}$ in Theorem 2 to obtain an alternative expression. Let $\Lambda^{\mathbf{y}_{n-1}}(x)=\left(\Lambda_{n}(x-y_{n-1})+\sum_{i=1}^{n-1}\Lambda_{i}(y_{i}-y_{i-1})\right)\wedge 1$ for $n\geqslant 2$, where $\mathbf{y}_{n-1}=(y_{1},\dots,y_{n-1})$ and $y_{0}=0$.
Proposition 4.
For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf_{\mathbf{y}_{n-1}\in\mathbb{R}^{n-1}}\Lambda^{\mathbf{y}_{n-1}}\mathrm{VaR}^{\mathbb{P}}(X).$
3.3 Conditional beliefs
In this subsection, we study further how the relationship among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ can affect the inf-convolution. In particular, we consider a case where agents operate with different information. To model this scenario, consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$. If $\mathbb{P}$ represents the physical probability, and $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are the beliefs of the agents, then those $\mathbb{Q}_{i}$ demonstrate that the agents’ may have different assessments of the occurrence of events due to the asymmetric information or being subject to different regulating agencies with heterogeneous stressing events; see e.g., Acciaio and Svindland (2009). For $B_{1},\dots,B_{n}\in\mathcal{F}$ and $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, let $\Lambda^{\diamond}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right)$ for all $x\in\mathbb{R}$. Note that $\Lambda^{\diamond}(x)$ can be viewed as the largest weighted aggregate risk tolerance of the agents to control the tail event $\{X>x\}$.
Theorem 3.
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~{}i=1,\dots,n$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, then
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=-\infty;$
If $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$ and $\Lambda^{\diamond}$ is attainable, then
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X),$
where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$. Moreover, if additionally $x^{*}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X)>-\infty$ and $\Lambda^{\diamond}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}=\Lambda^{\diamond}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}^{*},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{*}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})$.
Note that Theorem 3 shows that the optimal sets $(A_{1}^{*},\dots,A_{n}^{*})$ may not be unique. Under the assumption of Theorem 3, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}=\Lambda^{\diamond}(x^{*})$, optimal sets can be chosen as
$A_{i}^{*}$ =$=\left\{1-\sum_{j=1}^{i}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{i-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\},~{}i=1,\dots,n-1,$
$A_{n}^{*}$ =$=\left\{\left(1-\sum_{j=1}^{n}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right)\wedge 0<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\}$
with the convention that $\sum_{j=1}^{0}a_{j}=0$ and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$. Moreover, the right-continuity of $\Lambda^{\diamond}$ at $x^{*}$ can be removed if all $\Lambda_{i}$ are decreasing. Now, if $\Lambda^{\diamond}(x^{*})<1$, the probability of the optimal sets, under the probability $\mathbb{Q}$ representing evaluations with respect to the information set that all agents agree on, are given by
$\mathbb{Q}(A_{i}^{*})=\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})},\quad i=1,\dots,n-1,~{}\mathbb{Q}(A_{n}^{*})=1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})},$
or, equivalently,
$\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\cap A_{i}^{*}\right)=\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*}),~{}i=1,\dots,n-1,~{}\mathbb{P}(A_{n}^{*})=\mathbb{P}(\cap_{i=1}^{n}B_{i})-\sum_{j=1}^{n-1}\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*}).$
In this special case, the interplay among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ is characterized by the relationship between the sets $B_{1},\dots,B_{n}$. Following the conclusion in Proposition 3, those sets $B_{1},\dots,B_{n}$ appear in the expressions of the inf-convolution and the optimal allocation, suggesting the influence of the relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ on the inf-convolution. In particular, if the agents have a strong disagreement on the interested events such that $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, it leads to a trivial case. Note that the crucial difference to the homogeneous case is that here, in characterising the allocations, we use the set $\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{*}$ which includes both the set $A_{i}^{*}$ on which the contingent loss is paid and the conditioning set (information) on which all agents can agree. If all Lambda functions in Proposition 3 are assumed to be constants, then we immediately derive the explicit formulas for the inf-convolution of multiple $\mathrm{VaR}$ under conditional beliefs.
Corollary 1.
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~{}i=1,\dots,n$. For $\alpha_{i}\in(0,1)$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, then
$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=-\infty;$
If $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, then
$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{1\wedge\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}}^{\mathbb{Q}}(X),$
where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$.
The results in Corollary 1 extend the elegant formula for the inf-convolution of $\mathrm{VaR}$ under homogeneous beliefs to the case with conditional beliefs; see Corollary 2 of Embrechts et al. (2018). The simplicity of the expressions for $\mathrm{VaR}$ allow some additional insight. We can see that the trivial outcome $\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=-\infty$ is not only achieved in the case of full disagreement between agents $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$. It also occurs under weaker disagreement, when $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\geqslant 1$. This inequality may be satisfied, when $\mathbb{P}(\cap_{i=1}^{n}B_{i})$ is sufficiently small, indicating high – but partial – disagreement in conditional beliefs, and the levels $\alpha_{i}$ are sufficiently large, indicating a high level of risk tolerance.
It is worth mentioning that the results in Proposition 3 can be explained from the perspective of Co-risk measures. Recently, Co-risk measures are introduced in the context of systemic risk measurement, with reference to conditional $\mathrm{VaR}$ (Adrian and Brunnermeier, 2016; Girardi and Tolga Ergün, 2013) and conditional $\mathrm{ES}$ (Mainik and Schaanning, 2014) . Let $B\in\mathcal{F}$ satisfying $\mathbb{P}(B)>0$. For $\alpha\in(0,1)$, the conditional $\mathrm{VaR}$ $(\mathrm{Co\mathrm{VaR}})$ and conditional $\mathrm{ES}$ $(\mathrm{Co\mathrm{ES}})$ are defined respectively as
$\mathrm{VaR}_{\alpha}^{\mathbb{P}}(X|B)$ $=\inf\{x:\mathbb{P}(X\leqslant x|B)\geqslant 1-\alpha\},~{}\mathrm{ES}_{\alpha}^{\mathbb{P}}(X|B)=\frac{1}{\alpha}\int_{0}^{\alpha}\mathrm{VaR}_{t}^{\mathbb{P}}(X|B)\mathrm{d}t.$
Note that if $B=\Omega$, conditional $\mathrm{VaR}$ and $\mathrm{ES}$ boil down to the classical $\mathrm{VaR}$ and $\mathrm{ES}$, i.e., $\mathrm{VaR}_{\alpha}^{\mathbb{P}}(X)=\mathrm{VaR}_{\alpha}^{\mathbb{P}}(X|\Omega)$ and $\mathrm{ES}_{\alpha}^{\mathbb{P}}(X)=\mathrm{ES}_{\alpha}^{\mathbb{P}}(X|\Omega)$. Analogously, the conditional $\Lambda\mathrm{VaR}$ ($\mathrm{Co\Lambda\mathrm{VaR}}$) can be defined as
$\Lambda\mathrm{VaR}^{\mathbb{P}}(X|B)=\inf\{x:\mathbb{P}(X\leqslant x|B)\geqslant\Lambda(x)\}.$
Note that $\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X|B_{i})$ if $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$. Hence, the results in Theorem 3 can be interpreted as the risk sharing using multiple $\mathrm{Co\Lambda\mathrm{VaR}}$ to represent the preference of the agents. In the regulatory context, those $B_{i}$ could be stressing test events given by different supervising agencies.
3.4 Absolutely continuous beliefs
Next, we consider a different relation among the probability measures, absolute continuity. Due to the complexity of the problem, we here only consider the case $n=2$.
Theorem 4.
Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ and denote $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ by $\eta$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
$\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:g_{x}((1-F_{X}^{\mathbb{Q}_{1}}(x)-\Lambda_{1}(y))_{+})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R}\right\},$
where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t})=t$. Moreover, if $x^{*}=\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)>-\infty$ and $g_{x^{*}}((1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+})\leqslant\Lambda_{2}(x^{*}-y^{*})$ for some $y^{*}\in\mathbb{R}$, then the optimal allocation is given by
$X_{1}=\left(X-x^{*}\right)\mathds{1}_{\Omega\setminus A^{*}}+y^{*},~{}X_{2}=\left(X-x^{*}\right)\mathds{1}_{A^{*}}+x^{*}-y^{*},$ (4)
where $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}$ with $z^{*}$ satisfying $\mathbb{Q}_{1}(X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*})=(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}$.
In Theorem 4, we do not obtain explicit formulas as the previous propositions. This is due to the complexity of the relation between $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, which can be observed from the expression of $g_{x}$ depending on the joint distribution of $(X,\eta)$ under $\mathbb{Q}_{1}$. Let us next see some examples of $g_{x}$ and the corresponding $A^{*}$ in (4):
  1. (i)
    If $X$ and $\eta$ are comonotonic, then $g_{x}(t)=\int_{F_{X}^{\mathbb{Q}_{1}}(x)}^{F_{X}^{\mathbb{Q}_{1}}(x)+t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$;
  2. (ii)
    If $X$ and $\eta$ are countermonotonic, then $g_{x}(t)=\int_{0}^{t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\}$;
  3. (iii)
    If $X$ and $\eta$ are independent under $\mathbb{Q}_{1}$, then $g_{x}(t)=(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{t/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ with the convention that $/0=0$ 0 and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant\frac{(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}}{1-F_{X}^{\mathbb{Q}_{1}}(x^{*})}\}$ with $U_{\eta}^{\mathbb{Q}_{1}}$ being independent of $X$ under $\mathbb{Q}_{1}$.
Note that the optimal set $A^{*}$ in (i) and (ii) have very different structures. To demonstrate this, suppose both $X$ and $\eta$ are continuous random variables under $\mathbb{Q}_{1}$. In (i), $A^{*}$ can be rewritten as $A^{*}=\{F_{X}^{\mathbb{Q}_{1}}(x^{*})<U_{X}^{\mathbb{Q}_{1}}<(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$. Hence, the contingent loss of agent 2 takes place on a set that does not include the full right tail of $X$. This is understandable because the comonotonicity of $X$ and $\eta$ means that agent 2 considers the distribution of aggregate loss more risky in first-order stochastic dominance, $F_{X}^{\mathbb{Q}_{2}}\succeq_{\mathrm{st}}F_{X}^{\mathbb{Q}_{1}}$ (see e.g. Pesenti et al. (2019)). In the alternative case (ii) of comonotonicity, we can write $A^{*}=\{U_{X}^{\mathbb{Q}_{1}}>F_{X}^{\mathbb{Q}_{1}}(x^{*})+\Lambda_{1}(y^{*})\}$, such that agent 2 is exposed to the full tail of $X$, consistent with $F_{X}^{\mathbb{Q}_{1}}\succeq_{\mathrm{st}}F_{X}^{\mathbb{Q}_{2}}$. Hence, by (4), we see that if $X$ and $\eta$ are comonotonic, a lower contingent loss is allocated to the second agent; if $X$ and $\eta$ are countermonotonic, a higher contingent loss is allocated to the second agent.
Note that Theorem 4 can be simplified when the Lambda functions are constants.
Corollary 2.
Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ and denote $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ by $\eta$. For $\alpha_{1},\alpha_{2}\in(0,1)$, we have
$\mathrm{VaR}_{\alpha_{1}}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{VaR}_{\alpha_{2}}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:g_{x}((1-F_{X}^{\mathbb{Q}_{1}}(x)-\alpha_{1})_{+})\leqslant\alpha_{2}\right\},$
where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t})=t$.
Under the assumption of Corollary 2, we have the following results.
Corollary 3.
  1. (i)
    If $X$ and $\eta$ are comonotonic, then
    $\mathrm{VaR}_{\alpha_{1}}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{VaR}_{\alpha_{2}}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:\int_{F_{X}^{\mathbb{Q}_{1}}(x)}^{F_{X}^{\mathbb{Q}_{1}}(x)\vee(1-\alpha_{1})}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s\leqslant\alpha_{2}\right\};$
  2. (ii)
    If $X$ and $\eta$ are countermonotonic, then
    $\mathrm{VaR}_{\alpha_{1}}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{VaR}_{\alpha_{2}}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:\int_{0}^{(1-F_{X}^{\mathbb{Q}_{1}}(x)-\alpha_{1})_{+}}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s\leqslant\alpha_{2}\right\};$
  3. (iii)
    If $X$ and $\eta$ are independent under $\mathbb{Q}_{1}$, then
    $\mathrm{VaR}_{\alpha_{1}}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{VaR}_{\alpha_{2}}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{(1-F_{X}^{\mathbb{Q}_{1}}(x)-\alpha_{1})_{+}/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s\leqslant\alpha_{2}\right\}$
    with the convention that $/0=0$ 0.
4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogenous beliefs
Throughout this section, we set $\mathcal{X}=L^{\infty}$. We investigate the risk sharing problem between two agents, considering the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and a monotone risk measure $\rho^{\mathbb{Q}_{2}}$, where $\mathbb{Q}_{2}$ may be different from $\mathbb{Q}_{1}$. Note that $\rho$ here is very general including many commonly used risk measures as special cases.
Theorem 5.
Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\rho$ is monotone. Then we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$ (5)
Moreover, the existence of the optimal allocation for the inf-convolution is equivalent to the existence of the minimizer of (5). If $(x^{*},y^{*},B^{*})$ is the minimizer, then one optimal allocation is given by
$X_{1}^{*}=(X-x^{*}-y^{*})\mathds{1}_{(B^{*})^{c}}+x^{*},~{}X_{2}^{*}=(X-x^{*}-y^{*})\mathds{1}_{B^{*}}+y^{*}.$ (6)
The optimal allocation in (6) has the same form as that of Theorem 1: each agent is allocated a constant risk and a contingent risk, which is the excess of the total risk over the aggregate cash allocation restricted to some contingent event. Note that the expression of the inf-convolution in Theorem 5 involves an optimization problem with three parameters $(x,y,B)$, where $B$ is a measurable set contained in $\mathcal{F}$. The set $B$ makes this optimization problem complicated. The optimal allocation in (6) may not be pairwise countermonotonic.
Our conclusion in Theorem 5 extends the results of Embrechts et al. (2020) for $n=2$, where the inf-convolution of $\mathrm{VaR}$ or $\mathrm{ES}$ with heterogeneous beliefs is considered. If $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then Theorem 5 coincides with the results in Theorem 3 of Liu (2024). We notice that for the scenario that $\mathbb{Q}_{1}=\mathbb{Q}_{2}$ and $\Lambda$ is a constant, Theorem 5 extends the results in Theorem 2 of Liu et al. (2022) ($\rho$ is there considered to be a monetary tail risk measure) and Theorem 5.3 of Wang and Wei (2018) ($\rho$ is there a distortion risk measure). The result of Theorem 5 is still new even if $\Lambda\mathrm{VaR}$ is reduced to $\mathrm{VaR}$. Let us now consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$. Under this setup, the contingent set $B$ in (5) is given explicitly.
Proposition 5.
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~{}i=1,2$. Moreover, suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\rho$ is monotone. Then we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\},$ (7)
where $A_{x}=\emptyset$ if $\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})$ and $A_{x}=(B_{1}\cap B_{2}^{c})\cup(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})$ if $\mathbb{P}(B_{1}\cap B_{2}^{c})<(1-\Lambda(x))\mathbb{P}(B_{1})$ with $\alpha_{x}$ satisfying $\mathbb{P}(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})=(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Moreover, the existence of the optimal allocation of the inf-convolution is equivalent to the existence of the minimizer of (7). If $(x^{*},y^{*})$ is the minimizer, then one optimal allocation is given by
$X_{1}^{*}=(X-x^{*}-y^{*})\mathds{1}_{A_{x^{*}}^{c}}+x^{*},~{}X_{2}^{*}=(X-x^{*}-y^{*})\mathds{1}_{A_{x^{*}}}+y^{*}.$ (8)
Note that in the optimal allocation given by (8), the contingent event $A_{x^{*}}$ is a fixed event, making the optimal allocation more concrete than that of Theorem 5. We next interpret the optimal allocation given by (8) by focusing on the case $A_{x^{*}}=\emptyset$. Note that the condition for $A_{x^{*}}=\emptyset$ can be reformulated as the condition:
$\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x^{*}))\mathbb{P}(B_{1})\Leftrightarrow\mathbb{Q}_{1}(B_{2})\leqslant\Lambda(x^{*}).$
Clearly, $A_{x^{*}}=\emptyset$ means that all the risk stays with agent 1. The condition $\mathbb{Q}_{1}(B_{2})\leqslant\Lambda(x^{*})$ is satisfied when $\mathbb{Q}_{1}(B_{2})$ is small (this means that agent 1 dismisses the view of agent 2, so we have more information asymmetry); or when $\Lambda(x^{*})$ is large (this means that agent 1 has a higher risk tolerance.)
Let $\mathcal{G}=\{g:[0,1]\to[0,1]|~{}g~{}\text{is increasing and left-continuous satisfying}~{}g(0)=0~{}\text{and}~{}g(1)=1\}$. For $g\in\mathcal{H}$, the distortion risk measure $\rho_{g}^{\mathbb{Q}}$ is defined as
$\rho_{g}^{\mathbb{Q}}(X)$ $=\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
Distortion risk measures form a popular class of risk measures applied in insurance pricing, performance evaluation, decision theory and many other topics; see e.g., Föllmer and Schied (2016), Wang (1996), Wang (2000), Cherny and Madan (2009) and Yaari (1987). Note that if $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|B)$ for some $B\in\mathcal{F}$ satisfying $\mathbb{P}(B)>0$, then $\rho_{g}^{\mathbb{Q}}(X)$ corresponds to the Co-distortion risk measures introduced in Dhaene et al. (2022).
Furthermore, for an increasing function $u$, the expected utility is defined as $\mathbb{E}_{u}^{\mathbb{Q}}:X\mapsto\mathbb{E}^{\mathbb{Q}}(u(X))$.
In the following corollary, let $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{1}\cap B_{2})>0$, indicating that beliefs are not totally incompatible, and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|B_{1}\cap B_{2})$, which in a sense represents the consensus view of the two agents. Applying Proposition 5, we obtain the inf-convolution of $\mathrm{\Lambda\mathrm{VaR}}$ and $\rho_{g}/\mathbb{E}_{u}/\Lambda_{1}\mathrm{VaR}^{+}$ under heterogeneous beliefs, respectively. In order to simplify the expression, we suppose $\lambda^{+}$ is attainable for $\Lambda$, i.e., there is $x\in\mathbb{R}$ such that $\Lambda(x)=\lambda^{+}$.
Corollary 4.
Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\lambda^{+}$ is attainable. Then the following conclusion holds with $\beta=\frac{(1-\lambda^{+})\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$.
  1. (i)
    For $g\in\mathcal{G}$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$ then
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=\left\{\begin{array}[]{cc}\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t),&g(\beta)=1\\ -\infty,&g(\beta)<1\end{array}\right.;$
  2. (ii)
    For an increasing function $u$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\left(x+(1-\beta_{x})u(-\infty)+\int_{0}^{\beta_{x}}u\left(\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)-x\right)\mathrm{d}t\right),$
    where $u(-\infty)=\lim_{y\to-\infty}u(y)$ and $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$;
  3. (iii)
    For $\Lambda_{1}\in\mathcal{H}$ with $10<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<$ 1, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\overline{\Lambda}^{x}\mathrm{VaR}^{+,\mathbb{Q}}(X),$
    where $\overline{\Lambda}^{x}(z)=\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\wedge 1$.
Note that if we let $g(t)=(t/\alpha)\wedge 1$ for $\alpha\in(0,1)$, then the distortion risk measure is the Expected Shortfall, i.e., $\rho_{g}=\mathrm{ES}_{\alpha}$. In that case, part (i) of Corollary 4 simplifies to stating that, if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\left\{\begin{array}[]{cc}\frac{1}{\alpha}\int_{0}^{\alpha}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}t,&\alpha\leqslant\beta\\ -\infty,&\alpha>\beta\end{array}\right..$
For all three cases in Corollary 4 the condition $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, indicating high levels of belief heterogeneity / high level of risk tolerance for agent 1, gives a sufficient condition for the inf-convolution to be infinity. For part (i), we also get an infinite allocation as long as the quantity $\beta$ is small enough, such that $g(\beta)<1$. Note that in the special case where $\rho_{g}$ is Expected Shortfall at level $\alpha$, this means $\beta<\alpha$. To interpret small values of $\beta$ we can reformulate this quantity as follows:
$\beta=\frac{\mathbb{P}(B_{1})}{\mathbb{P}(B_{2})}\left(\mathbb{Q}_{1}(B_{2})-\lambda^{+}\right).$
It is seen that $\beta$ becomes small if
  • $\lambda^{+}$ is high, such that agent 1 has a high maximal risk tolerance;
  • $\mathbb{Q}_{1}(B_{2})$ is low, so we have more belief heterogeneity;
  • $\frac{\mathbb{P}(B_{1})}{\mathbb{P}(B_{2})}$ is low, since we are in the case that $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. This fraction can be understood, in some sense, as a comparison of the subjective informativeness of the conditioning done by the two agents. If a low probability event is observed, one could claim that more has been learned about the underlying risk. Hence, if the ratio is small, agent 1 has subjectively learned more than agent 2 and is, in that specific sense, more confident.
These interpretations tell us that we have an infinite allocation when there is belief heterogeneity and agent 1 is more risk tolerant as well as considering that the event they believe they observed is highly informative.
We mention that the results in Corollary 4 extend the results in the literature in the sense that $\Lambda$ is not a constant and the second risk measure is not law-invariant under $\mathbb{Q}_{1}$. For instance, (i) of Corollary 4 exactly extends Theorem 5.3 of Wang and Wei (2018) ($\rho$ is a distortion risk measure which is law-invariant under $\mathbb{Q}_{1}$) in these two directions. It is worth mentioning that (iii) of Corollary 4 is a new result even for the case $\mathbb{Q}_{1}=\mathbb{Q}_{2}$.
Let us now consider the case $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$. Instead of finding a simplified version of Theorem 5 for this case, we apply Theorem 5 directly to some specific $\rho$. For $\alpha\in(0,1)$, let $\mathrm{LES}_{\alpha}^{\mathbb{Q}}(X)=\frac{1}{\alpha}\int_{1-\alpha}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t$, which represents the lower conditional tail expectation.
Corollary 5.
Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ with $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ denoted by $\eta$. For $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1, we have the following conclusions.
  1. (i)
    For $\alpha\in(0,1)$,
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\inf_{t\in\mathbb{R}}\left(t+\frac{1-\lambda^{+}}{\alpha}\mathrm{LES}_{1-\lambda^{+}}^{\mathbb{Q}_{1}}(\eta(X-t)_{+})\right);$
  2. (ii)
    For an increasing function $u$, we have
    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\left(x+u(-\infty)\Lambda(x)\mathrm{ES}_{\Lambda(x)}^{\mathbb{Q}_{1}}(\eta)+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})\right),$
    where $u(-\infty)=\lim_{y\to-\infty}u(y)$.
Note that in (i) of Corollary 5, if $\Lambda$ is a constant, then the result coincides with Theorem 5 of Embrechts et al. (2020) for $n=2$, where the semi-explicit formula is offered. For the case $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$, it might be difficult to find more explicit expression for the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and $\rho_{g}^{\mathbb{Q}_{2}}$/$\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$.
5 Inf-convolution of $\Lambda\mathrm{VaR}^{+}$ and a monotone risk measure
The inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is a very challenging problem even if we suppose the beliefs are homogeneous; see the discussion for homogeneous beliefs in Liu (2024). One important feature of $\Lambda\mathrm{VaR}^{+}$ is to capture the tail risk for decreasing $\Lambda$; see Frittelli et al. (2014) and Hitaj et al. (2018). In this section, we focus on the inf-convolution of $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}$ and a monotone risk measure $\rho^{\mathbb{Q}_{2}}$, which includes the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ as a special case. Let $\mathcal{X}=L^{\infty}$. For any $\Lambda\in\mathcal{H}$, we denote $\Lambda_{x}(z)=\inf_{x\leqslant t\leqslant z}\Lambda(t),~{}z\geqslant x$. Then $\Lambda_{x}(z)$ is decreasing and right-continuous for $z\geqslant x$. If $\Lambda$ is decreasing and right-continuous, then $\Lambda_{x}(z)=\Lambda(z),~{}~{}z\geqslant x$; and if $\Lambda$ is increasing, then $\Lambda_{x}(z)=\Lambda(x),~{}~{}z\geqslant x$. The following result is valid for general $\Lambda\in\mathcal{H}$.
Theorem 6.
Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1, and $\rho$ is monotone and law-invariant under $\mathbb{Q}_{2}$. Then we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\},$ (9)
where
$\Lambda_{x,y}(z)=\left\{\begin{array}[]{cc}0,&z<x\\ 1-\Lambda_{x}(z),&x\leqslant z<y\\ 1,&z\geqslant y\end{array}\right..$
Moreover, the optimal allocation of the inf-convolution exists if and only if the minimizer of (9) exists. If $(x^{*},y^{*},U^{*})$ is the minimizer, then one optimal allocation is given by
$X_{1}^{*}=\Lambda_{x^{*},y^{*}}^{-1}(U^{*}),~{}X_{2}^{*}=X-\Lambda_{x^{*},y^{*}}^{-1}(U^{*}).$ (10)
In the case of $\rho=\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$ for $\Lambda_{1}\in\mathcal{H}$ with $10<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<$ 1, it follows from Theorem 6 that the optimal risk allocation for risk sharing between $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}$ and $\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$ has the form of (10), which is neither comonotonic nor pairwise countermonotonic. This means that the optimal risk sharing for multiple $\Lambda\mathrm{VaR}^{+,\mathbb{Q}}$ is a very different problem from the cases with convex risk measures or quantile-based risk measures in the literature. The expression of the inf-convolution in (9) is an optimization problem involving three parameters $(x,y,U)$, where $U\sim U[0,1]$ under $\mathbb{Q}_{1}$. Note that the distribution of $U$ under $\mathbb{Q}_{2}$ depends on the relation between $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$. Moreover, the expression in (9) and also the optimal allocation in (10) show that shape of $\Lambda$ function plays an important role in both the expression of the inf-convolution and the form of the optimal allocations, which is very different from the results of $\Lambda\mathrm{VaR}$ displayed in Sections 3 and 4.
We are also aware that the conclusion in Theorem 6 is complicated in the sense that it involves $\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x,y}^{-1}(U))$, which is a very difficult problem even if $\mathbb{Q}_{1}=\mathbb{Q}_{2}$. For $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, $\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x,y}^{-1}(U))$ corresponds to the problem of robust risk aggregation with dependence uncertainty, i.e., finding the value of $\inf_{X\sim F,Y\sim G}\rho(X+Y)$, where the marginal distributions are known but the dependence structure is completely unknown. We refer to Wang and Wang (2016), Jakobsons et al. (2016), Blanchet et al. (2024), Han and Liu (2024) and the references therein for the discussion on the robust risk aggregation.
In order to simplify (9), we consider a special case of $\Lambda$ functions. This makes it easier to find more explicit expressions of the inf-convolution for concrete examples of $\rho$.
Proposition 6.
For $x_{1}\in\mathbb{R}$ and $10<\lambda_{1}<\lambda_{2}<$ 1, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$. Under the assumption of Theorem 6, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-x\mathds{1}_{B_{1}}-x\vee x_{1}\mathds{1}_{B_{2}}-y\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)\right\},$ (11)
where $\mathcal{B}_{\lambda_{1},\lambda_{2}}=\{(B_{1},B_{2}):\mathbb{Q}_{1}(B_{1})=\lambda_{1},\mathbb{Q}_{2}(B_{2})=\lambda_{2}-\lambda_{1},B_{1}\cap B_{2}=\emptyset\}$.
Note that the optimization problem in (11) is much easier than that in (9) because it only has parameters $(x,y,B_{1},B_{2})$.
6 Conclusion
In this paper, we consider the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under heterogenous beliefs. We obtain the general expression for the inf-convolution and the corresponding optimal allocations. To study impact of the relation of the beliefs on the inf-convolution, we discuss three different cases: i) homogeneous beliefs; ii) conditional beliefs; iii) absolute continuous beliefs. For all these cases, we obtain more explicit expressions of the inf-convolution. In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high. The inf-convolution of one ΛVaR/$\Lambda\mathrm{VaR}/\Lambda\mathrm{VaR}^{+}$ and a general risk measures with belief heterogeneity are also discussed. There are still some unsolved problems such as the inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs, which deserves further study.
Acknowledgements: YW acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2023-04674).
7 Declarations
Conflict of interest: All authors declare that they have no conflict of interest.
Appendix A Proof of Section 2
Proof of Proposition 1. Note that there exists $A\in\mathcal{F}$ such that $\mathbb{Q}_{1}(A)=0$ and $\mathbb{Q}_{2}(A)=1$. It follows from the law-invariance of $\rho_{1}$ and $\rho_{2}$ and the constant preservation of $\rho_{2}$ that
$\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)$ $\leqslant\rho_{1}^{\mathbb{Q}_{1}}(X-c\mathds{1}_{A})+\rho_{2}^{\mathbb{Q}_{2}}(c\mathds{1}_{A})$
$=\rho_{1}^{\mathbb{Q}_{1}}(X)+\rho_{2}^{\mathbb{Q}_{2}}(c)=\rho_{1}^{\mathbb{Q}_{1}}(X)+c\to-\infty$
as $c\to-\infty$. Hence, $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)=-\infty.$ This completes the proof. ∎
Appendix B Proof of Section 3
In this section, we give all of the proofs of the results in Section 3.
Proof of Theorem 1. We first show that $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. For $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, let $x=\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$ and $y_{i}=\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$. Note that $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 implies $y_{i}\in\mathbb{R}$ and $x\in\mathbb{R}$. By the definition of $\Lambda_{i}\mathrm{VaR}$ and the right continuity of $\Lambda_{i}$, we have $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$ with $C_{i}=\{X_{i}>y_{i}\}$. We denote $D_{i}=C_{i}\cup\left(\cup_{i=1}^{n}C_{i}\right)^{c}$. Then it follows that
$\{X>x\}=\left\{\sum_{i=1}^{n}X_{i}>\sum_{i=1}^{n}y_{i}\right\}\subset\bigcup_{i=1}^{n}C_{i}.$
For $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, let $A_{i}\subset D_{i}$. Then we have
$\{X>x,A_{i}\}\subset\{X>x,D_{i}\}=\{X>x,C_{i}\}\subset C_{i}.$
Using the fact $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$, we have $\mathbb{Q}_{i}(X>x,A_{i})\leqslant\Lambda(y_{i})$. This implies $\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})=x\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. By the arbitrary of $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, we have $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
Next we show the inverse inequality. Suppose there exist $y_{i}\in\mathbb{R},(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$. Let $X_{i}=(X-x)\mathds{1}_{A_{i}}+y_{i},~{}i=1,\dots,n$, where $x=\sum_{i=1}^{n}y_{i}$. Then $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$. Direct computation gives
$\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})=y_{i}+\Lambda_{i}^{y_{i}}\mathrm{VaR}^{\mathbb{Q}_{i}}((X-x)\mathds{1}_{A_{i}}),$
where $\Lambda_{i}^{y_{i}}(z)=\Lambda_{i}(z+y_{i})$ for $z\in\mathbb{R}$. Note that $\mathbb{Q}_{i}((X-x)\mathds{1}_{A_{i}}>0)=\mathbb{Q}_{i}(X>x,A_{i})\leqslant\Lambda_{i}(y_{i})=\Lambda_{i}^{y_{i}}(0)$. By definition, we have $\Lambda_{i}^{y_{i}}\mathrm{VaR}^{\mathbb{Q}_{i}}((X-x)\mathds{1}_{A_{i}})\leqslant 0$. Hence,
$\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})\leqslant\sum_{i=1}^{n}y_{i}=x.$
This implies $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\leqslant\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})\leqslant\sum_{i=1}^{n}y_{i}$. By the arbitrary of $y_{i}$, we have $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\leqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. One can directly check that $X_{i}^{*},~{}i=1,\dots,n$ is the optimal allocation. We complete the proof. ∎
Proof of Proposition 2. Case (i). Note that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\leqslant 1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ implies $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{-}$ for all $i=1,\dots,n$. Hence, for any $(y_{1},\dots,y_{n})$, it follows that $\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$. Consequently, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$.
Case (ii). Let $x^{*}=\bigwedge_{i=1}^{n}\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}X$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ represents the $\mathrm{ess\mbox{-}inf}$ under $\mathbb{Q}_{i}$. For any $(y_{1},\dots,y_{n})$ with $\sum_{i=1}^{n}y_{i}<x^{*}$ and $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, there exists $i\in\{1,\dots,n\}$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}^{n}y_{i},A_{i})=\mathbb{Q}_{i}(A_{i})>\lambda_{i}^{+}\geqslant\Lambda_{i}(y_{i})$. Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)\geqslant x^{*}>-\infty$. ∎
Proof of Proposition 3. We first consider the case that $\Lambda_{i}\in\mathcal{H}$ is decreasing. If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{+}$ for all $i=1,\dots,n$. Note that $\lim_{y_{i}\to-\infty}\Lambda(y_{i})=\lambda_{i}^{+}$. Hence, there exists $y_{i}^{0}$ such that $\Lambda(y_{i})\geqslant\mathbb{Q}_{i}(A_{i})$ for all $y_{i}<y_{i}^{0}$. It follows that $\mathbb{Q}_{i}(X>\sum_{i=1}^{n}y_{i},A_{i})\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\Lambda_{i}(y_{i})$ if $y_{i}\leqslant y_{i}^{0}$ for all $i=1,\dots,n$. This implies $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$. The second statement is shown in (ii) of Proposition 2.
Next, we consider the case that $\Lambda_{i}\in\mathcal{H}$ is increasing. If $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then there exists $i\in\{1,\dots,n\}$ such that $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{-}$ and $\mathbb{Q}_{j}(A_{j})<\lambda_{j}^{+}$ for $j\neq i$. Note that there exists $y_{j}^{0}\in\mathbb{R}$ such that $\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j}^{0})$ for all $j\neq i$. Moreover, for any $y_{i}\in\mathbb{R}$, it follows that $\mathbb{Q}_{j}(X>y_{i}+\sum_{j\neq i}y_{j}^{0},A_{j})\leqslant\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j})$ for $j\neq i$, and $\mathbb{Q}_{i}(X>y_{i}+\sum_{j\neq i}y_{j}^{0},A_{i})\leqslant\Lambda_{i}(y_{i})$. By the arbitrary of $y_{i}$, we have $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$.
We next show the last statement. We denote $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)=1+\varepsilon$ for some $\varepsilon>0$. Suppose by contradiction that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$. Then there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ satisfying $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=-\infty$, and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}y_{i}^{(m)},A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$. Without loss of generality, we can assume $\lim_{m\to\infty}y_{i}^{(m)}=y_{i}$. Let $E_{1}=\{i\in\{1,\dots,n\}:y_{i}=-\infty\}$ and $E_{2}=\{i\in\{1,\dots,n\}:y_{i}>-\infty\}$. By the fact that $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=-\infty$, we have $E_{1}\neq\emptyset$. Moreover, there exists $m_{0}>0$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}y_{i}^{(m)})=1$ for all $m\geqslant m_{0}$. Hence, we have $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$ for all $m\geqslant m_{0}$. Note that $\lim_{m\to\infty}\Lambda_{i}(y_{i}^{(m)})=\lambda_{i}^{-}$ for all $i\in E_{1}$. Hence, for any $0<\eta<\varepsilon$, there exists $m_{1}\geqslant m_{0}$ such that $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{-}(1+\eta)$ for all $m\geqslant m_{1}$ and $i\in E_{1}$. Moreover, $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{+}$ for all $m\geqslant m_{1}$ and $i\in E_{2}$. Combination of the those conclusion leads to $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i}^{(m)})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)\leqslant 1+\eta<1+\varepsilon$ for $m>m_{1}$, which is a contradiction of the assumption. Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$. We complete the proof. ∎
Proof of Theorem 2. By Theorem 1, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Note that $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ if and only if $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})$. Hence, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)$ $=\inf\left\{\sum_{i=1}^{n}y_{i}:(y_{1},\dots,y_{n})\in\mathbb{R}^{n},\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right\}$
$=\inf\left\{x\in\mathbb{R}:\sum_{i=1}^{n}y_{i}=x,\mathbb{P}\left(X>x\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right\}$ (12)
$=\inf\left\{x\in\mathbb{R}:\mathbb{P}\left(X>x\right)\leqslant\Lambda^{*}(x)\right\}=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X).$
For the optimal allocation, note that the right continuity of $\Lambda^{*}$ at $x^{*}$ implies $\mathbb{P}\left(X>x^{*}\right)\leqslant\Lambda^{*}(x^{*})$. Moreover, there exist $(y_{1}^{*},\dots,y_{n}^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ such that $1\bigwedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $\mathbb{P}\left(X>x^{*},A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$. Hence, $(y_{1}^{*},\dots,y_{n}^{*},A_{1}^{*},\dots,A_{n}^{*})$ is the minimizer of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. Using Theorem 1, the claimed allocation is an optimal allocation. ∎
Proof of Proposition 4. In light of (12), we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)$ $=\inf\left\{x\in\mathbb{R}:\sum_{i=1}^{n}y_{i}=x,\mathbb{P}\left(X>x\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right\}$
$=\inf_{\mathbf{y}_{n-1}\in\mathbb{R}^{n-1}}\Lambda^{\mathbf{y}_{n-1}}\mathrm{VaR}^{\mathbb{P}}(X).$
Proof of Theorem 3. By Theorem 1, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$
$=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Let $\mathcal{N}_{n}=\{\{i_{1},\dots,i_{m}\}\subset\{1,\dots,n\}:~{}m=1,2,\dots,n-1\}$. For $\{i_{1},\dots,i_{m}\}\in\mathcal{N}_{n}$, we denote $C_{\{i_{1},\dots,i_{m}\}}=\left(\bigcap_{i\in\{i_{1},\dots,i_{m}\}}B_{i}\right)\setminus\left(\bigcup_{i\in\{1,\dots,n\}\setminus\{i_{1},\dots,i_{m}\}}B_{i}\right)$. Then for $\{i_{1},\dots,i_{m}\},\{i_{1}^{\prime},\dots,i_{m^{\prime}}^{\prime}\}\in\mathcal{N}_{n}$, we have $C_{\{i_{1},\dots,i_{m}\}}\cap C_{\{i_{1}^{\prime},\dots,i_{m^{\prime}}^{\prime}\}}=\emptyset$ if $\{i_{1},\dots,i_{m}\}\neq\{i_{1}^{\prime},\dots,i_{m^{\prime}}^{\prime}\}$. Moreover, it follows that $\bigcup_{\{i_{1},\dots,i_{m}\}\in\mathcal{N}_{n}}C_{\{i_{1},\dots,i_{m}\}}=\left(\bigcup_{i=1}^{n}B_{i}\right)\setminus\left(\bigcap_{i=1}^{n}B_{i}\right)$. Next, for each $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, we do some operations on it. We fix $\{i_{1},\dots,i_{m}\}\in\mathcal{N}_{n}$. For each $i\in\{i_{1},\dots,i_{m}\}$, we do the following operations:
$A_{i}\to A_{i}\setminus C_{\{i_{1},\dots,i_{m}\}},~{}A_{j}\to A_{j}\cup\left(A_{i}\cap C_{\{i_{1},\dots,i_{m}\}}\right),$
where $j=\min(\{1,\dots,n\}\setminus\{i_{1},\dots,i_{m}\})$. The new sets after the operations are denoted by $(A_{1}^{\prime},\dots,A_{n}^{\prime})$. Clearly, $(A_{1}^{\prime},\dots,A_{n}^{\prime})\in\Pi_{n}(\Omega)$ and $B_{i}\cap A_{i}^{\prime}\subset B_{i}\cap A_{i}$ for all $i=1,\dots,n$ and $A_{i}^{\prime}\cap C_{\{i_{1},\dots,i_{m}\}}=\emptyset$ for all $i\in\{i_{1},\dots,i_{m}\}$. Those operations will be done for all $\{i_{1},\dots,i_{m}\}\in\mathcal{N}_{n}$. The final sets are denoted by $(A_{1}^{\prime\prime},\dots,A_{n}^{\prime\prime})$. Clearly, $(A_{1}^{\prime\prime},\dots,A_{n}^{\prime\prime})\in\Pi_{n}(\Omega)$ and $B_{i}\cap A_{i}^{\prime\prime}\subset B_{i}\cap A_{i}$. Moreover, it follows that $A_{i}^{\prime\prime}\cap C_{\{i_{1},\dots,i_{m}\}}=\emptyset$ for all $\{i_{1},\dots,i_{m}\}\in\mathcal{N}_{n}$ satisfying $i\in\{i_{1},\dots,i_{m}\}$. This implies $B_{i}\cap A_{i}^{\prime\prime}=\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{\prime\prime}$ for all $i=1,\dots,n$. By the construction of $A_{i}^{\prime\prime}$, it follows that $\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\subset B_{i}\cap A_{i}^{\prime\prime}$, which together with the fact $B_{i}\cap A_{i}^{\prime\prime}\subset B_{i}\cap A_{i}$ implies $\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}=\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{\prime\prime}=B_{i}\cap A_{i}^{\prime\prime}$. Hence, we have
$\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}$
$\leqslant\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Note that the inverse inequality is trivial. Hence, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$
$=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Clearly, $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=-\infty$ if $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)=0$. Next, we consider the case $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)>0$. Note that $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ and some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ is equivalent to $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\cap_{i=1}^{n}B_{i}\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$. Using this conclusion and the attainability of $\Lambda^{\diamond}$, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$ $=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{P}\left(X>x,\cap_{i=1}^{n}B_{i}\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})\right\}$
$=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{Q}\left(X>x\right)\leqslant\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\}$
$=\inf\left\{x:\mathbb{Q}\left(X>x\right)\leqslant\Lambda^{\diamond}(x)\right\}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X).$
One can easily check that the claimed allocation is optimal. ∎
Proof of Theorem 4. Using Theorem 1, we have
$\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)$
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{Q}_{2}\left(X>x,A\right)\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},A\in\mathcal{F}\right\}$
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A\}})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},A\in\mathcal{F}\right\}.$
Let $f(t)=\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant t),~{}t\in[0,1]$. Clearly, $f$ is an increasing and continuous function. Let $f^{-1}$ denote the left-quantile of $f$. Define $A_{x,t}=\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant f^{-1}(t)\}$ for $t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$. Let $t_{0}=\mathbb{Q}_{1}(X>x,A)$. Then it follows that $\mathbb{Q}_{1}(X>x,A_{x,t_{0}})=\mathbb{Q}_{1}(X>x,A)$ and
$\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A_{x,t_{0}}\}})=\mathbb{E}^{\mathbb{Q}_{1}}(F_{\eta}^{\mathbb{Q}_{1},-1}(U_{\eta}^{\mathbb{Q}_{1}})\mathds{1}_{\{X>x,A_{x,t_{0}}\}})\leqslant\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A\}}).$
Consequently, we have
$\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)$
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A_{x,t}^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A_{x,t}\}})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
$=\inf\left\{x\in\mathbb{R}:1-F_{X}^{\mathbb{Q}_{1}}(x)-t\leqslant\Lambda_{1}(y),g_{x}(t)\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
$=\inf\left\{x\in\mathbb{R}:g_{x}((1-F_{X}^{\mathbb{Q}_{1}}(x)-\Lambda_{1}(y))_{+})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R}\right\}.$
Next, we show $(X_{1},X_{2})$ given by (4) is the optimal allocation. Note that
$\mathbb{Q}_{1}(X_{1}>y^{*})$ $=\mathbb{Q}_{1}(X>x^{*},\Omega\setminus A^{*})=\mathbb{Q}_{1}(X>x^{*})-\mathbb{Q}_{1}(X>x^{*},A^{*})$
$=1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\leqslant\Lambda_{1}(y^{*}),$
$\mathbb{Q}_{2}(X_{2}>x^{*}-y^{*})$ $=\mathbb{Q}_{2}(X>x^{*},A^{*})=\mathbb{E}^{\mathbb{Q}_{1}}\left(\eta\mathds{1}_{\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}}\right)$
$=g_{x^{*}}((1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+})\leqslant\Lambda_{2}(x^{*}-y^{*}).$
Hence, we have $\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})\leqslant y^{*}$ and $\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X_{2})\leqslant x^{*}-y^{*}$, which implies $\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X_{2})\leqslant\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)$. By the definition of the inf-convolution, $(X_{1},X_{2})$ is the optimal allocation. This completes the proof. ∎
Appendix C Proof of Section 4
In this section, we provide all the proofs of results in Section 4.
Proof of Theorem 5. For any $x,y$ and $B$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, let $X_{1}=x\mathds{1}_{B}+(X-y)\mathds{1}_{B^{c}}$. Direct calculation gives
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})\leqslant x+\rho^{\mathbb{Q}_{2}}(X-X_{1})=x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}).$
This implies $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)\leqslant x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})$. By the arbitrary of $x,y$ and $B$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)\leqslant\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
We next show the inverse inequality. For any $X_{1}\in\mathcal{X}$, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$. Note that $X_{1}=F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ a.s. under $\mathbb{Q}_{1}$. Define $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ and $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{i}}$ represent the $\mathrm{ess\mbox{-}inf}$ and $\mathrm{ess\mbox{-}sup}$ under $\mathbb{Q}_{i}$, $i=1,2$. Note that $U_{X_{1}}^{\mathbb{Q}_{1}}$ can be chosen such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$. By the definition of $\Lambda\mathrm{VaR}$ and the right continuity of $\Lambda$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x_{1})\geqslant 1-\Lambda(x_{1})$, which implies $x_{1}\geqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(1-\Lambda(x_{1}))$. Hence we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(Y_{1})=x_{1}$, and $X_{1}\leqslant Y_{1}$ for all $w\in\Omega$. By monotonicity of $\rho$, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})$ $\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-Y_{1})$
$=x_{1}+\rho^{\mathbb{Q}_{2}}((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}})$
$\geqslant\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
By the arbitrary of $X_{1}$, we obtain
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)\geqslant\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
Combing the above conclusions, we have (5) holds.
We next show that the existence of the optimal allocation implies the existence of the minimizer of (5). Suppose there exists $X_{1}\in\mathcal{X}$ such that $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$. Following the same argument as above to show the inverse inequality, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$ and $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$ such that $X_{1}\leqslant Y_{1}$ for all $w\in\Omega$. We have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})$ $\geqslant\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(Y_{1})+\rho^{\mathbb{Q}_{2}}(X-Y_{1})$
$=x_{1}+\rho^{\mathbb{Q}_{2}}((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}})$
$\geqslant\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
Using $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$ and (5), we have
$x_{1}+\rho^{\mathbb{Q}_{2}}((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}})=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
This implies $(x_{1},-m,\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\})$ is the minimizer of (5).
Now suppose $(x^{*},y^{*},B^{*})$ is the minimizer of (5). We next check $(X_{1}^{*},X_{2}^{*})$ given in (6) is an optimal allocation. It follows that
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1}^{*})+\rho^{\mathbb{Q}_{2}}(X_{2}^{*})$ $\leqslant x^{*}+\rho^{\mathbb{Q}_{2}}((X-x^{*})\mathds{1}_{B^{*}}+y^{*}\mathds{1}_{(B^{*})^{c}})$
$=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}$
$=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X).$
Consequently, $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1}^{*})+\rho^{\mathbb{Q}_{2}}(X_{2}^{*})=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$ and $(X_{1}^{*},X_{2}^{*})$ is an optimal allocation. This completes the proof. ∎
Proof of Proposition 5. In light of Theorem 5, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
If $\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})$, then there exists $D\in\mathcal{F}$ such that $D\subseteq B_{1}\cap B_{2}^{c}$ and $\mathbb{Q}_{1}(D)=1-\Lambda(x)$. Using the law-invariance of $\rho$ under $\mathbb{Q}_{2}$, we have
$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}$ $\leqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{D}+y\mathds{1}_{D^{c}})\right\}$
$=\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
For $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have $(X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\geqslant y$ a.s. under $\mathbb{Q}_{2}$. It follows from the monotonicity and law-invariance of $\rho$ that $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant\rho^{\mathbb{Q}_{2}}(y)$. Using the monotonicity of $\rho$ again, we have $\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$ Consequently,
$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}=\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
Next, we consider the case $\mathbb{P}(B_{1}\cap B_{2}^{c})<(1-\Lambda(x))\mathbb{P}(B_{1})$. Let $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$ and $A_{x}=(B_{1}\cap B_{2}^{c})\cup(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})$ with $\alpha_{x}$ satisfying $\mathbb{P}(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})=(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. For any $B\in\mathcal{F}$ with $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $z\geqslant y$, we have
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\frac{\mathbb{P}(X\leqslant z+x,B\cap B_{2})+\mathbb{P}(B^{c}\cap B_{2})}{\mathbb{P}(B_{2})}=1-\frac{\mathbb{P}(X>z+x,B\cap B_{2})}{\mathbb{P}(B_{2})}.$
Observe that
$\mathbb{P}(X>z+x,A_{x}\cap B_{2})=\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})\leqslant\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap B_{3})$
for any $B_{3}\in\mathcal{F}$ such that $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Note that $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ implies $\mathbb{P}(B_{1}\cap B_{2}\cap B)\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Hence we have $\mathbb{P}(X>z+x,A_{x}\cap B_{2})\leqslant\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap B)\leqslant\mathbb{P}(X>z+x,B\cap B_{2})$. This implies
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)\leqslant\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)$
for $z\geqslant y$. For $z<y$, clearly,
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)=0.$
By the law-invariance and monotonicity of $\rho$, we have for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}),$ which implies
$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}=\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\}.$
Hence, (7) holds. The conclusion on optimal allocation follows from the same reasoning as the proof of Theorem 5. The details are omitted. We complete the proof. ∎
Proof of Corollary 4. We first consider (i). Note that
$x+\rho_{g}^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})=\rho_{g}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}).$
If $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then there exists $x\in\mathbb{R}$ such that $\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})$. Hence, $A_{x}=\emptyset$. By Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(x+y)=-\infty$.
Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. Suppose $y<\inf^{\mathbb{Q}_{2}}X-x$. For $z\geqslant x+y$, we have
$\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)$ $=\frac{\mathbb{P}(A_{x}^{c}\cap B_{2})+\mathbb{P}(X\leqslant z,A_{x}\cap B_{2})}{\mathbb{P}(B_{2})}$ (13)
$=1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})},$
where $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$. Moreover, $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)=0$ for $z<x+y$. If $g(\beta)<1$, then there exists $x\in\mathbb{R}$ such that $g(\beta_{x})<1$. By (13), we have $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)\geqslant 1-\beta_{x}$ for all $z\geqslant x+y$, which implies $\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\leqslant x+y\to-\infty$ as $y\to-\infty$ for all $t\in[\beta_{x},1)$. Hence, for $x\in\mathbb{R}$ with $\beta_{x}=\beta$, we have
$\inf_{y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})=\lim_{y\to-\infty}\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)=-\infty.$
Now, we assume $g(\beta)=1$. Direct computation shows for $0<t<\beta_{x}$,
$\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})$ $=\inf\{z\geqslant x+y:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}\geqslant 1-t\}$
$=\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X).$ (14)
Hence, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\int_{[0,\beta)}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)$
$=\inf_{x\in\mathbb{R}}\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t)=\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
Let us now consider (ii). If $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then there exists $x\in\mathbb{R}$ such that $\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})$. Hence, $A_{x}=\emptyset$. By Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+u(y))=-\infty$.
Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. It follows from Proposition 5 and (C) that
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{2}}u((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\left(x+(1-\beta_{x})u(-\infty)+\int_{0}^{\beta_{x}}u\left(\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)-x\right)\mathrm{d}t\right),$
where $u(-\infty)=\lim_{y\to-\infty}u(y)$.
Finally, we consider (iii). Similarly as above, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then by Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+y)=-\infty$.
Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. In light of Proposition 5 and (13), we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)$
$=\inf_{x\in\mathbb{R}}\lim_{y\to-\infty}\left\{x+\sup\{z\geqslant y:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant x+z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z)\}\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\to-\infty}\sup\left\{z\geqslant y+x:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<\beta_{x}-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{1}\cap B_{2})}<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:F_{X}^{\mathbb{Q}}(z)<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}=\inf_{x\in\mathbb{R}}\overline{\Lambda}^{x}\mathrm{VaR}^{+,\mathbb{Q}}(X),$
where $\overline{\Lambda}^{x}(z)=\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\bigwedge 1$. We complete the proof. ∎
Proof of Corollary 5. We first consider (i). In light of Theorem 5, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X\mathds{1}_{B}+(x+y)\mathds{1}_{B^{c}})\right\}.$
It shows in Rockafellar and Uryasev (2002) that $\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\inf_{t\in\mathbb{R}}\left(t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}\left((X-t)_{+}\right)\right)$. Hence, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X\mathds{1}_{B}+(x+y)\mathds{1}_{B^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\inf_{t\in\mathbb{R}}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}((X-t)_{+}\mathds{1}_{B}+(x+y-t)_{+}\mathds{1}_{B^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{t\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{1}}(\eta(X-t)_{+}\mathds{1}_{B})\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{t\in\mathbb{R}}\left\{t+\frac{1}{\alpha}\int_{\Lambda(x)}^{1}\mathrm{VaR}_{s}^{\mathbb{Q}_{1}}(\eta(X-t)_{+})\mathrm{d}s\right\}$
$=\inf_{t\in\mathbb{R}}\left\{t+\frac{1}{\alpha}\int_{\lambda^{+}}^{1}\mathrm{VaR}_{s}^{\mathbb{Q}_{1}}(\eta(X-t)_{+})\mathrm{d}s\right\}.$
Next, we consider (ii). It follows from Theorem 5 that
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{2}}\left(u(X-x)\mathds{1}_{B}+u(y)\mathds{1}_{B^{c}}\right)\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{2}}\left(u(X-x)\mathds{1}_{B}+u(-\infty)\mathds{1}_{B^{c}}\right)\right\}.$
Note that $u(-\infty)\leqslant u(X-x)$. Hence, we have
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$
$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{B})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})+u(-\infty)\int_{0}^{\Lambda(x)}\mathrm{VaR}_{t}^{\mathbb{Q}_{1}}(\eta)\mathrm{d}t\right\}.$
Appendix D Proof of Section 5
In this section, we offer all the proofs of the results in Section 5.
Proof of Theorem 6. For $x\in\mathbb{R},y\geqslant x$ and $U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]$, we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(\Lambda_{x,y}^{-1}(U))=x$. Hence,
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho(X)\leqslant\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(\Lambda_{x,y}^{-1}(U))+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x,y}^{-1}(U))=x+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x,y}^{-1}(U)).$
Using the arbitrary of $x\in\mathbb{R},y\geqslant x$ and $U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]$, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)\leqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
Next, we show the inverse inequality. For $X_{1}\in\mathcal{X}$, we let $x_{1}=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$. We choose $U_{X_{1}}^{\mathbb{Q}_{1}}$ such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$. This implies $\{F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant x\}=\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant F_{X_{1}}^{\mathbb{Q}_{1}}(x)\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(F_{X_{1}}^{\mathbb{Q}_{1}}(x))\}\subset\{X_{1}\leqslant x\}$ for all $x\in\mathbb{R}$. Hence, $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ for all $w\in\Omega$. By definition, $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant 1-\Lambda(x)$ for all $x\geqslant x_{1}$. This implies $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for $x_{1}\leqslant x\leqslant y$. If $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, then $F_{X_{1}}^{\mathbb{Q}_{1}}(x)=1\geqslant\Lambda_{x_{1},y}(x)$ for $x\geqslant y$. Hence, if $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for all $x\in\mathbb{R}$. Combing all the above results, we have $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant\Lambda_{x_{1},y}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ if $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$. It follows that for $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$,
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})$ $\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))$
$\geqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
By the arbitrary of $X_{1}$, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)\geqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
Hence, we obtain (9).
Next, we show the optimal allocation of the inf-convolution exists if and only if the minimizer of (9) exists. Suppose $(X_{1},X-X_{1})$ is the optimal allocation of the inf-convolution. Then we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\square\rho^{\mathbb{Q}_{2}}(X)$. Let $x_{1}=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$ and $y_{1}>x_{1}\vee\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$. We choose $U_{X_{1}}^{\mathbb{Q}_{1}}$ such that $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ for all $w\in\Omega$. Using the above argument, we have $F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$. By the monotonicity of $\rho$, we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))$. This implies $x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}$. Hence, $(x_{1},y_{1},U_{X_{1}}^{\mathbb{Q}_{1}})$ is the minimizer of (9). Moreover, if $(x_{1},y_{1},U)$ is the minimizer of (9), one can easily check that $(\Lambda_{x_{1},y_{1}}^{-1}(U),X-\Lambda_{x_{1},y_{1}}^{-1}(U))$ is the optimal allocation of the inf-convolution. We complete the proof. ∎
Proof of Proposition 6. Direct calculation shows that if $x<x_{1}$, then $\Lambda_{x,y}(z)=\lambda_{1}\mathds{1}_{\{x\leqslant z<x_{1}\}}+\lambda_{2}\mathds{1}_{\{x_{1}\leqslant z<y\}}+\mathds{1}_{\{z\geqslant y\}}$; if $x\geqslant x_{1}$, then $\Lambda_{x,y}(z)=\lambda_{2}\mathds{1}_{\{x\leqslant z<y\}}+\mathds{1}_{\{z\geqslant y\}}$. Hence, we have $\Lambda_{x,y}^{-1}(t)=x\mathds{1}_{\{0<t\leqslant\lambda_{1}\}}+x\vee x_{1}\mathds{1}_{\{\lambda_{1}<t\leqslant\lambda_{2}\}}+y\mathds{1}_{\{\lambda_{2}<t<1\}}$. For any $U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]$, it follows that $\Lambda_{x,y}^{-1}(U)=x\mathds{1}_{\{0<U\leqslant\lambda_{1}\}}+x\vee x_{1}\mathds{1}_{\{\lambda_{1}<U\leqslant\lambda_{2}\}}+y\mathds{1}_{\{\lambda_{2}<U<1\}}$. Note that $\mathbb{Q}_{1}(0<U\leqslant\lambda_{1})=\lambda_{1}$, $\mathbb{Q}_{1}(\lambda_{1}<U\leqslant\lambda_{2})=\lambda_{2}-\lambda_{1}$ and $\mathbb{Q}_{1}(\lambda_{2}<U<1)=1-\lambda_{2}$, and those three sets are disjoint. Consequently, in light of Theorem 6, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-x\mathds{1}_{B_{1}}-x\vee x_{1}\mathds{1}_{B_{2}}-y\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)\right\}.$

### v3

@@
-
+Risk sharing with Lambda Value-at-Risk under heterogeneous beliefs
 Abstract
@@
 We obtain semi-explicit formulas for the inf-convolution of multiple Lambda Value-at-Risk measures under heterogeneous beliefs and the explicit forms of the corresponding optimal allocations.
-To show the impact of belief heterogeneity, we consider three cases: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs.
+To show the impact of belief heterogeneity, we consider three cases: homogeneous beliefs, conditional beliefs, and general beliefs with two agents.
 For those cases, we find more explicit expressions for the inf-convolution, showing the influence of the relation of the beliefs on the inf-convolution.
@@
 The relation between the beliefs plays a crucial role in the risk sharing problem; see e.g., Chi (2019).
-To study the impact of belief heterogeneity, we consider three cases in Section 3: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs.
+To study the impact of belief heterogeneity, we consider three cases in Section 3: homogeneous beliefs, conditional beliefs, and general beliefs with two agents.
 For homogeneous beliefs, we show that the inf-convolution of $\Lambda\mathrm{VaR}$ is still a $\Lambda\mathrm{VaR}$ for general $\Lambda$ functions, extending the results in Liu (2025) and Xia and Hu (2024), where the monotonicity of the $\Lambda$ functions is typically required.
@@
 Under this setup, we also obtain an explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$, showing that the inf-convolution of $\Lambda\mathrm{VaR}$ under conditional beliefs is also a $\Lambda\mathrm{VaR}$ under a new conditional belief.
-Conditional beliefs are closely related to the conditional risk measures proposed in the literature to measure systemic risk, such as $\mathrm{CoVaR}$ (Adrian and Brunnermeier (2016) and Girardi and Tolga Ergün (2013)) and $\mathrm{CoES}$ (Mainik and Schaanning (2014)).
+Conditional beliefs are closely related to the conditional risk measures proposed in the literature to quantify systemic risk, such as $\mathrm{CoVaR}$ (Adrian and Brunnermeier (2016) and Girardi and Tolga Ergün (2013)) and $\mathrm{CoES}$ (Mainik and Schaanning (2014)).
 The risk sharing problem in this case can be interpreted as risk sharing with some Co-risk measures ($\mathrm{Co\Lambda\mathrm{VaR}}$, a new conditional risk measure proposed in Section 3).
-For the third case, we consider the risk sharing between two agents and suppose one probability measure is absolutely continuous with another one.
-Under this assumption, we find a more explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$.
-Note that we also give the results for the inf-convolution of $\mathrm{VaR}$ with conditional beliefs and absolutely continuous beliefs respectively, which are also new to the literature.
-In Section 4, we study the inf-convolution of one $\Lambda\mathrm{VaR}$ and a general monotone risk measure with belief heterogeneity.
-The expression of the inf-convolution and the form of the optimal allocation are derived.
-Then we consider two cases: conditional beliefs and absolutely continuous beliefs.
-For conditional beliefs, we obtain explicit expressions for the inf-convolution of one $\Lambda\mathrm{VaR}$ and one distortion risk measure/expected utility/$\Lambda\mathrm{VaR}^{+}$.
-For absolutely continuous beliefs, we find the formula for the inf-convolution of one $\Lambda\mathrm{VaR}$ and one $\mathrm{ES}$/expected utility.
-In Section 5, we consider the inf-convolution of one $\Lambda\mathrm{VaR}^{+}$ (an alternative definition of Lambda Value-at-Risk) and a general monotone risk measure with belief heterogeneity.
-We obtain the expression of the inf-convolution, which is complicated, and the forms of the optimal allocation.
-Our results include the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs as a special case.
-The results here are not explicit and very different from the previous sections, as they involve the shape of the $\Lambda$ function and the best case of risk aggregation under dependence uncertainty.
-This means that $\Lambda\mathrm{VaR}^{+}$ is very different from $\Lambda\mathrm{VaR}$ for its application within the risk sharing problem.
+For the third case, we consider the risk sharing between two agents.
+With the aid of the Lebesgue decomposition theorem, we find a more explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$ with general heterogeneous beliefs.
+Note that we also give the explicit expression for the inf-convolution of $\mathrm{VaR}$ with heterogeneous beliefs, which are also new to the literature.
+[…]
 Throughout the paper, the stated results provide conditions for trivial, i.e., infinite, inf-convolutions.
@@
 Hence, our contribution illuminates how the entanglement of beliefs and preferences impacts key features of the risk sharing problem with $\Lambda\mathrm{VaR}$.
-The notation and key definitions are given in Section 2 and all the proofs are delegated to Appendix A-D.
+The notation and key definitions are given in Section 2 and all the proofs are delegated to Appendix A.
 2 Notation and Definitions
-For a given atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $L^{\infty}$ denote the collection of all bounded random variables.
+For a given probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $L^{\infty}$ denote the collection of all bounded random variables.
 Moreover, let $\mathcal{X}$ be a set of random variables containing $L^{\infty}$.
+We say $\mathcal{X}$ is unbounded if $\mathcal{X}$ contains unbounded random variables.
 We suppose $\mathcal{X}$ has good enough properties to conduct our study, such as $\mathcal{X}=L^{p}$ for some $p\geqslant 0$, where $L^{p}$ represents the set of all random variables with finite $p$-th moments.
-Throughout this paper, we assume that the probability measures $\mathbb{Q},\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ live on $(\Omega,\mathcal{F})$ and the corresponding probability spaces are atomless.
+Throughout this paper, we consider the random variables themselves but not the corresponding equivalent classes, because, as will be seen in the sequel, more than one probability measures are considered simultaneously.
 For any $X\in\mathcal{X}$, a positive value of $X$ represents a financial loss.
-The distribution function of $X$ under $\mathbb{Q}$ is denoted as $F_{X}^{\mathbb{Q}}$.
-For a mapping $\rho:\mathcal{X}\to\mathbb{R}$, we say $\rho$ is law-invariant under $\mathbb{Q}$ if for all $X,Y\in\mathcal{X}$,
+For a risk measure, that is, a mapping $\rho:\mathcal{X}\to\mathbb{R}$, we say that $\rho$ monotone if $X\leqslant Y$ implies $\rho(X)\leqslant\rho(Y)$; and $\rho$ is cash-additive if $\rho(X+c)=\rho(X)+c$ for $X\in\mathcal{X}$ and $c\in\mathbb{R}$.
+We say $\rho$ is a monetary risk measure if it is monotone and cash-additive.
+Furthermore, $\rho$ is law-invariant under a probability measure $\mathbb{Q}$ if for all $X,Y\in\mathcal{X}$,
 $X\overset{\mathbb{Q}}{=}Y~~\Rightarrow~~\rho(X)=\rho(Y),$ (1)
-where $\overset{\mathbb{Q}}{=}$ stands for equality in distribution under $\mathbb{Q}$; $\rho$ is monotone if $X\leqslant Y$ implies $\rho(X)\leqslant\rho(Y)$; and $\rho$ is cash-additive if $\rho(X+c)=\rho(X)+c$ for $X\in\mathcal{X}$ and $c\in\mathbb{R}$.
-We say $\rho$ is a monetary risk measure if it is monotone and cash-additive.
-For more details on risk measures, one can refer to Chapter 4 of Föllmer and Schied (2016).
-For simplicity, throughout the paper, we use the notation $\rho^{\mathbb{Q}}$ to represent that $\rho$ is law-invariant under $\mathbb{Q}$.
+[…]
+For a probability measure $\mathbb{Q}$, the distribution function of $X$ under $\mathbb{Q}$ is denoted as $F_{X}^{\mathbb{Q}}$.
 For $X\in\mathcal{X}$, $F_{X}^{\mathbb{Q},-1}$ represents its left-quantile under $\mathbb{Q}$, which is defined by
@@
 with the convention that $\inf\emptyset=\infty$.
-For any $X\in\mathcal{X}$, we denote by $U_{X}^{\mathbb{Q}}$ a uniform random variable on $[0,1]$ under $\mathbb{Q}$ such that $X=F_{X}^{\mathbb{Q},-1}(U_{X}^{\mathbb{Q}})$ a.s.
+For any $X\in\mathcal{X}$, if $\mathbb{Q}$ is atomless, we denote by $U_{X}^{\mathbb{Q}}$ a uniform random variable on $[0,1]$ under $\mathbb{Q}$ such that $X=F_{X}^{\mathbb{Q},-1}(U_{X}^{\mathbb{Q}})$ a.s.
 under $\mathbb{Q}$.
 The existence of such $U_{X}^{\mathbb{Q}}$ for any random variable $X$ is guaranteed by e.g., Lemma A.32 of Föllmer and Schied (2016).
+We next define tail risk measures, which are important for our results later. For a random variable $X\in\mathcal{X}$, $p\in(0,1]$ and an atomless probability measure $\mathbb{Q}$, we call
+$X_{p}=(F_{X}^{\mathbb{Q}})^{-1}(1-p+pU_{X}^{\mathbb{Q}})$
+the tail risk of $X$ beyond its (1p)(1-p)-quantile under $\mathbb{Q}$. The distribution of $X_{p}$ under $\mathbb{Q}$ is given by
+$\mathbb{Q}(X_{p}\leqslant x)=\frac{(F_{X}^{\mathbb{Q}}(x)-(1-p))_{+}}{p},~x\in\mathbb{R},$
+where $x_{+}=\max(x,0)$.
+We say $\rho$ is a $p$-tail risk measure under $\mathbb{Q}$ for some $p\in(0,1)$ if $X_{p}\overset{\mathbb{Q}}{=}Y_{p}$ implies $\rho^{\mathbb{Q}}(X)=\rho^{\mathbb{Q}}(Y)$ for all $X,Y\in\mathcal{X}$.
+One can refer to Liu and Wang (2021) and Liu et al.
+(2022) for more details on the definition, properties and applications of tail risk measures.
 Next, we define the inf-convolution. The random variable $X\in\mathcal{X}$ represents a risk to be shared between $n$ agents. Then, define the set of allocations of $X$ as
 $\mathbb{A}_{n}(X)=\left\{(X_{1},\ldots,X_{n})\in\mathcal{X}^{n}:\sum_{i=1}^{n}X_{i}=X\right\}.$
+Throughout this paper, we assume that the probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ live on $(\Omega,\mathcal{F})$ and, with the exception of Subsection 3.1, the corresponding probability spaces are atomless.
 In this paper, we suppose the agents may have heterogeneous beliefs, represented by some probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$.
-The inf-convolution of the risk measures $\rho_{1}^{\mathbb{Q}_{1}},\dots,\rho_{n}^{\mathbb{Q}_{n}}$ is the mapping $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}:\mathcal{X}\to[-\infty,\infty)$, defined as
+The inf-convolution of the risk measures $\rho_{1},\dots,\rho_{n}$, under the respective probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ is the mapping $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}:\mathcal{X}\to[-\infty,\infty)$, defined as
 $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)=\inf\left\{\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i}):(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)\right\}.$
 Note that $\rho_{i}$ is law-invariant under $\mathbb{Q}_{i}$, where $\mathbb{Q}_{i}$ represents the belief of the i-th agent.
-An $n$-tuple $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is called an optimal allocation of $X$ for $(\rho_{1}^{\mathbb{Q}_{1}},\dots,\rho_{n}^{\mathbb{Q}_{n}})$ if $\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i})=\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$.
-The inf-convolution $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}(X)$ can be interpreted as the smallest possible aggregate capital for the total risk $X$ in financial system, if $\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ represents the capital charge for the i-th financial institution to hold the risky position $X_{i}$.
+In case of belief heterogeneity, the inf-convolution may not be law-invariant, i.e., if $X\overset{\mathbb{Q}_{i}}{=}Y$ for all $i=1,\dots,n$, then $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)\neq\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(Y)$ may hold; see more details in the Remark at the end of Section 3.3.
+An $n$-tuple $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is called an optimal allocation of $X$ for $(\rho_{1},\dots,\rho_{n})$ under $(\mathbb{Q}_{1},\dots,\mathbb{Q}_{n})$ if $\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i})=\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$.
+The inf-convolution $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$ can be interpreted as the smallest possible aggregate capital for the total risk $X$ in financial system, if $\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ represents the capital charge for the i-th financial institution to hold the risky position $X_{i}$.
 More economic interpretations on the inf-convolution can be found in e.g., Delbaen (2012), Rüschendorf (2013) and Embrechts et al.
 (2018).
-For the risk measures $(\rho_{1},\dots,\rho_{n})$ and a total risk $X$, an allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is Pareto-optimal if for any other allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$, $\rho_{i}(Y_{i})\leqslant\rho_{i}(X_{i})$ for all $i=1,\dots,n$ implies $\rho_{i}(Y_{i})=\rho_{i}(X_{i})$ for all $i=1,\dots,n$.
+For the risk measures $(\rho_{1},\dots,\rho_{n})$ under $(\mathbb{Q}_{1},\dots,\mathbb{Q}_{n})$ and a total risk $X$, an allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is Pareto-optimal if for any other allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$, $\rho_{i}^{\mathbb{Q}_{i}}(Y_{i})\leqslant\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ for all $i=1,\dots,n$ implies $\rho_{i}^{\mathbb{Q}_{i}}(Y_{i})=\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ for all $i=1,\dots,n$.
 For finite-valued monetary risk measures, it is shown in Embrechts et al.
@@
 where $\inf\emptyset=\infty$ and $\sup\emptyset=$-$\infty$.
-Note that $\Lambda\mathrm{VaR}(X)=\Lambda\mathrm{VaR}^{+}(X)$ if $\Lambda$ is increasing; Otherwise, it may not be true; see Proposition 6 of Bellini and Peri (2022).
+Note that $\Lambda\mathrm{VaR}^{\mathbb{Q}}(X)=\Lambda\mathrm{VaR}^{+,\mathbb{Q}}(X)$ if $\Lambda$ is increasing; Otherwise, it may not be true; see Proposition 6 of Bellini and Peri (2022).
 We refer to Bellini and Peri (2022) for two other definitions of $\Lambda\mathrm{VaR}$.
@@
 This means that if the beliefs of agents are too divergent, then each one of them can get unreasonably good deal by transferring risks between them.
-Then, the fact that the inf-convolution is $\infty$- ∞ reflects that agents can always make their position better and hence there is no optimal allocation.
+Then, the fact that the inf-convolution is -$\infty$ reflects that agents can always make their position better and hence there is no optimal allocation.
 In contrast, if $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)>$-$\infty$ for many monetary risk measures $\rho_{1}$ and $\rho_{2}$; see e.g., Embrechts et al.
@@
 Suppose $\mathcal{X}\supseteq L^{\infty}$.
-We find an expression for the inf-convolution and also the corresponding optimal allocation, covering the results of the inf-convolution of $\Lambda\mathrm{VaR}$ or $\mathrm{VaR}$ with homogeneous or heterogenous beliefs in Liu (2024), Embrechts et al.
+We find an expression for the inf-convolution and also the corresponding optimal allocation, covering the results of the inf-convolution of $\Lambda\mathrm{VaR}$ or $\mathrm{VaR}$ with homogeneous or heterogeneous beliefs in Liu (2025), Embrechts et al.
 (2020) and Embrechts et al.
 (2018).
-We first consider the general beliefs and then discuss three different cases about the relation of the beliefs: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs.
+We first consider the general beliefs and then discuss three different cases about the relation of the beliefs: homogeneous beliefs, conditional beliefs, and general beliefs with two agents.
 3.1 General beliefs
-In this subsection, we suppose $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are general probability measures.
+In this subsection only, we assume that $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are general probability measures that are not required to be atomless.
 We first introduce the following notation.
-Let $\Pi_{n}(\Omega)=\{(A_{1},\dots,A_{n}):\cup_{i=1}^{n}A_{i}=\Omega,~{}A_{i}\cap A_{j}=\emptyset,~{}i\neq j\}$.
-For $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, let
-$\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
+Let $\Pi_{n}(\Omega)=\{(A_{1},\dots,A_{n}):\cup_{i=i}^{n}A_{i}=\Omega,~A_{i}\cap A_{j}=\emptyset,~i\neq j\}$.
+For $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, let
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{Q}_{i}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$}
 Note that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)\in[-\infty,\infty)$.
 Theorem 1.
-For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
+For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
 $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X).$
 Moreover, if the minimizer of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$ exists denoted by $(y_{1}^{*},\dots,y_{n}^{*},A_{1}^{*},\dots,A_{n}^{*})$, then the optimal allocation is given by
-$X_{i}=\left(X-\sum_{i=1}^{n}y_{i}^{*}\right)\mathds{1}_{A_{i}^{*}}+y_{i}^{*},~{}i=1,\dots,n.$ (3)
+$X_{i}=\left(X-\sum_{j=1}^{n}y_{j}^{*}\right)\mathds{1}_{A_{i}^{*}}+y_{i}^{*},~i=1,\dots,n.$ (3)
 The allocation (3) can be understood as follows.
@@
 Compared with the results in Liu (2025), our results in Theorem 1 offer semi-explicit formulas for the inf-convolution.
-Later, we will specify the relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ to obtain more explicit formulas for the inf-convolution.
-Moreover, Theorem 1 does not explicitly show when the inf-convolution is $\infty$- ∞.
+Later, we will specify some special relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ to obtain more explicit formulas for the inf-convolution.
+However, it is unclear what the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under general heterogeneous beliefs is, leaving it as an open problem and deserving further study in future.
+Moreover, Theorem 1 does not explicitly show when the inf-convolution is -$\infty$.
 The discussion on the finiteness of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$ is complicated as it heavily relies on the properties of Lambda functions and the relationship between $\mathbb{Q}_{i}$; see Liu (2025) and Proposition 1 in Section 2.
@@
 Proposition 2.
-Let $\mathcal{X}=L^{\infty}$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
+For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
   1. (i)
@@
   2. (ii)
-    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ for all $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
-Note that if all $\Lambda_{i}$ are constants, then Proposition 2 offers a sufficient and necessary condition for the finiteness of the inf-convolution.
-In general, due to the complexity of the $\Lambda_{i}$ functions, those conditions are sufficient but not necessary.
+    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
+Note that if all $\Lambda_{i}$ are constants and $\mathbb{Q}_{1}=\mathbb{Q}_{2}=\dots=\mathbb{Q}_{n}$, then Proposition 2 offers a sufficient and necessary condition for the finiteness of the inf-convolution.
+In general, due to the complexity of the $\Lambda_{i}$ functions and the heterogeneity of probability measures, those conditions are sufficient but not necessary.
 As we can see from Proposition 2, cases (i) and (ii) cannot cover all scenarios.
@@
 Proposition 3.
-Let $\mathcal{X}=L^{\infty}$. Suppose all $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 are decreasing. Then we have the following conclusion.
+Suppose all $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$ are decreasing. Then we have the following conclusion.
   1. (i)
-    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$;
+    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$;
   2. (ii)
-    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ for all $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
-Suppose all $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 are increasing. Then we have the following conclusion.
+    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
+Suppose all $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$ are increasing. Then we have the following conclusion.
   1. (i)
@@
 In this more refined result, we see that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ plays a pivotal role when $\Lambda$ functions are decreasing.
-We see that, in order to guarantee a fininte inf-convolution for decreasing $\Lambda_{i}$, it is necessary for this quantity to exceed 1.
+We see that, in order to guarantee a finite inf-convolution for decreasing $\Lambda_{i}$, it is necessary for this quantity to exceed 1.
 For simplicity, assume that for all $i$, $\lambda_{i}^{+}$ is attained by $\Lambda_{i}$.
@@
 This means that for each agent we can find a set $A_{i}$ on which the risk is suitably bounded by the largest acceptable probability to the agent of an adverse tail event.
-Motivated by this, we may consider $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ as a measure of belief homogeneity, which also reflects preferences.
+We can consider $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ as a measure of belief homogeneity, which also reflects preferences.
 Consider two extreme cases.
@@
 On the other hand, consider the case of belief homogeneity, $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$.
-Then, we will have $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ if and only if $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, which reflects the impact of preferences.
+If $\mathbb{P}$ is atomless, then we will have $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ if and only if $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, which reflects the impact of preferences.
 Clearly, when the values of $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, indicating that agents accept lower total probabilities of adverse events, the inf-convolution is finite.
@@
 3.2 Homogeneous beliefs
-In this subsection, we focus on the case $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$.
-We find more explicit expression for the inf-convolution.
-Let $\Lambda^{*}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right)$ for all $x\in\mathbb{R}$.
+In this subsection, we focus on the case $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. We find more explicit expression for the inf-convolution. For $x\in\mathbb{R}$, let
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda^{*}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right)$}.
 We say $\Lambda^{*}$ is attainable if for each $x\in\mathbb{R}$, there exists $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ satisfying $\sum_{i=1}^{n}y_{i}=x$ such that $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})=\Lambda^{*}(x)$.
@@
 Theorem 2.
-For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, if $\Lambda^{*}$ is attainable, then we have
+For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, if $\Lambda^{*}$ is attainable, then we have
 $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X).$
-Moreover, if additionally $x^{*}=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)>-\infty$ and $\Lambda^{*}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(X>x^{*},A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
+Moreover, if additionally $x^{*}=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)>$-$\infty$ and $\Lambda^{*}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(\{X>x^{*}\}\cap A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
 Note that Theorem 2 shows that the optimal sets $(A_{1}^{*},\dots,A_{n}^{*})$ may not be unique.
 Under the assumption of Theorem 2, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$, optimal sets can be chosen as
-$A_{i}^{*}$ =$=\left\{1-\sum_{j=1}^{i}\Lambda_{j}(y_{j}^{*})<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{i-1}\Lambda_{j}(y_{j}^{*})\right\},~{}i=1,\dots,n-1,$
-$A_{n}^{*}$ =$=\left\{\left(1-\sum_{j=1}^{n}\Lambda_{j}(y_{j}^{*})\right)\wedge 0<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{n-1}\Lambda_{j}(y_{j}^{*})\right\},$
+$A_{i}^{*}$ $=\left\{1-\sum_{j=1}^{i}\Lambda_{j}(y_{j}^{*})<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{i-1}\Lambda_{j}(y_{j}^{*})\right\},~i=1,\dots,n-1,$
+$A_{n}^{*}$ $=\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(1-\sum_{j=1}^{n}\Lambda_{j}(y_{j}^{*})\right)\vee 0}<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{n-1}\Lambda_{j}(y_{j}^{*})\right\},$
 with the convention that $\sum_{j=1}^{0}a_{i}=0$.
@@
 In particular, Theorem 2 covers the case that some $\Lambda\mathrm{VaR}$ have increasing $\Lambda$ functions and others have decreasing $\Lambda$ functions, demonstrating that the agents may have relatively different risk appetites.
-We can also remove the assumption on the attainability of $\Lambda^{*}$ in Theorem 2 to obtain an alternative expression.
-Let $\Lambda^{\mathbf{y}_{n-1}}(x)=\left(\Lambda_{n}(x-y_{n-1})+\sum_{i=1}^{n-1}\Lambda_{i}(y_{i}-y_{i-1})\right)\wedge 1$ for $n\geqslant 2$, where $\mathbf{y}_{n-1}=(y_{1},\dots,y_{n-1})$ and $y_{0}=0$.
+We can also remove the assumption on the attainability of $\Lambda^{*}$ in Theorem 2 to obtain an alternative expression. For $n\geqslant 2$, let
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda^{\mathbf{y}_{n-1}}(x)=\left(\Lambda_{n}(x-y_{n-1})+\sum_{i=1}^{n-1}\Lambda_{i}(y_{i}-y_{i-1})\right)\wedge 1$},
+where $\mathbf{y}_{n-1}=(y_{1},\dots,y_{n-1})$ and $y_{0}=0$.
 Proposition 4.
-For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
+For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
 $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf_{\mathbf{y}_{n-1}\in\mathbb{R}^{n-1}}\Lambda^{\mathbf{y}_{n-1}}\mathrm{VaR}^{\mathbb{P}}(X).$
+Different from Theorem 1, the assumption of measures being atomless cannot be removed in Theorem 2. This can be seen from the following example.
+Example 1.
+For $m\geqslant 2$, suppose there exist distinct $w_{1},w_{2},\dots,w_{m}\in\Omega$ such that $\{w_{1}\},\dots,\{w_{n}\}\in\mathcal{F}$ and $\mathbb{P}({w_{1}})=\dots=\mathbb{P}({w_{m}})=1/m$.
+Define a random variable $X$ with $X(w_{i})=i,~i=1,\dots,m$ and $X(w)=0$ if $w\notin\{w_{1},\dots,w_{m}\}$.
+Clearly, $X$ is measurable under $(\Omega,\mathcal{F})$.
+Moreover, suppose $\Lambda_{1}=\Lambda_{2}=1/(2m)$.
+It follows from Theorem 1 that
+$~~\Lambda_{1}\mathrm{VaR}^{\mathbb{P}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{P}}(X)$
+$=\inf\{y\in\mathbb{R}:\mathbb{P}(\{X>y\}\cap A))\vee\mathbb{P}(\{X>y\}\cap A^{c}))\leqslant 1/(2m)~\text{for some}~A\in\mathcal{F}\}=m.$
+On the other hand,
+$\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)=\mathrm{VaR}_{1/m}^{\mathbb{P}}(X)=m-1.$
+Hence, $\Lambda_{1}\mathrm{VaR}^{\mathbb{P}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{P}}(X)\neq\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)$.
+This implies that the conclusion in Theorem 2 may not hold if the atomlessness assumption is removed.
 3.3 Conditional beliefs
@@
 If $\mathbb{P}$ represents the physical probability, and $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are the beliefs of the agents, then those $\mathbb{Q}_{i}$ demonstrate that the agents’ may have different assessments of the occurrence of events due to the asymmetric information or being subject to different regulating agencies with heterogeneous stressing events; see e.g., Acciaio and Svindland (2009).
-For $B_{1},\dots,B_{n}\in\mathcal{F}$ and $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, let $\Lambda^{\diamond}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right)$ for all $x\in\mathbb{R}$.
-Note that $\Lambda^{\diamond}(x)$ can be viewed as the largest weighted aggregate risk tolerance of the agents to control the tail event $\{X>x\}$.
+For $B_{1},\dots,B_{n}\in\mathcal{F}$ and $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, let
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda^{\diamond}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}\right)$}
+for all $x\in\mathbb{R}$. Note that $\Lambda^{\diamond}(x)$ can be viewed as the largest weighted aggregate risk tolerance of the agents to control the tail event $\{X>x\}$.
 Theorem 3.
 Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~i=1,\dots,n$.
-For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, then
+For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, then
 $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=-\infty;$
@@
 where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$.
-Moreover, if additionally $x^{*}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X)>-\infty$ and $\Lambda^{\diamond}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}=\Lambda^{\diamond}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}^{*},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{*}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})$.
+Moreover, if additionally $x^{*}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X)>$-$\infty$ and $\Lambda^{\diamond}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}=\Lambda^{\diamond}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(\{X>\sum_{i=1}^{n}y_{i}^{*}\}\cap\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{j}^{*}\right)\leqslant\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})$.
 Note that Theorem 3 shows that the optimal sets $(A_{1}^{*},\dots,A_{n}^{*})$ may not be unique.
-Under the assumption of Theorem 3, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}=\Lambda^{\diamond}(x^{*})$, optimal sets can be chosen as
-$A_{i}^{*}$ =$=\left\{1-\sum_{j=1}^{i}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{i-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\},~{}i=1,\dots,n-1,$
-$A_{n}^{*}$ =$=\left\{\left(1-\sum_{j=1}^{n}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right)\wedge 0<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\}$
+Under the assumption of Theorem 3, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}=\Lambda^{\diamond}(x^{*})$, optimal sets can be chosen as
+$A_{i}^{*}$ $=\left\{1-\sum_{j=1}^{i}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{i-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}\right\},~i=1,\dots,n-1,$
+$A_{n}^{*}$ $=\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(1-\sum_{j=1}^{n}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}\right)\vee 0}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}\right\}$
 with the convention that $\sum_{j=1}^{0}a_{j}=0$ and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$.
@@
 Now, if $\Lambda^{\diamond}(x^{*})<1$, the probability of the optimal sets, under the probability $\mathbb{Q}$ representing evaluations with respect to the information set that all agents agree on, are given by
-$\mathbb{Q}(A_{i}^{*})=\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})},\quad i=1,\dots,n-1,~{}\mathbb{Q}(A_{n}^{*})=1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})},$
+$\mathbb{Q}(A_{i}^{*})=\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})},\quad i=1,\dots,n-1,~\mathbb{Q}(A_{n}^{*})=1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})},$
 or, equivalently,
-$\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\cap A_{i}^{*}\right)=\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*}),~{}i=1,\dots,n-1,~{}\mathbb{P}(A_{n}^{*})=\mathbb{P}(\cap_{i=1}^{n}B_{i})-\sum_{j=1}^{n-1}\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*}).$
+$\mathbb{P}\left((\cap_{k=1}^{n}B_{k})\cap A_{i}^{*}\right)=\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*}),~i=1,\dots,n-1,~\mathbb{P}(A_{n}^{*})=\mathbb{P}(\cap_{k=1}^{n}B_{k})-\sum_{j=1}^{n-1}\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*}).$
 In this special case, the interplay among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ is characterized by the relationship between the sets $B_{1},\dots,B_{n}$.
@@
 In particular, if the agents have a strong disagreement on the interested events such that $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, it leads to a trivial case.
-Note that the crucial difference to the homogeneous case is that here, in characterising the allocations, we use the set $\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{*}$ which includes both the set $A_{i}^{*}$ on which the contingent loss is paid and the conditioning set (information) on which all agents can agree.
+Note that the crucial difference to the homogeneous case is that here, in characterising the allocations, we use the set $\left(\cap_{k=1}^{n}B_{k}\right)\cap A_{i}^{*}$ which includes both the set $A_{i}^{*}$ on which the contingent loss is paid and the conditioning set (information) on which all agents can agree.
 If all Lambda functions in Proposition 3 are assumed to be constants, then we immediately derive the explicit formulas for the inf-convolution of multiple $\mathrm{VaR}$ under conditional beliefs.
@@
 If $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, then
-$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{1\wedge\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}}^{\mathbb{Q}}(X),$
+$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{1\wedge\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}}^{\mathbb{Q}}(X),$
 where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$.
@@
 We can see that the trivial outcome $\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=$-$\infty$ is not only achieved in the case of full disagreement between agents $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$.
-It also occurs under weaker disagreement, when $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\geqslant 1$.
-This inequality may be satisfied, when $\mathbb{P}(\cap_{i=1}^{n}B_{i})$ is sufficiently small, indicating high – but partial – disagreement in conditional beliefs, and the levels $\alpha_{i}$ are sufficiently large, indicating a high level of risk tolerance.
+It also occurs under weaker disagreement, when $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}\geqslant 1$.
+This inequality may be satisfied, when $\mathbb{P}(\cap_{j=1}^{n}B_{j})$ is sufficiently small, indicating high – but partial – disagreement in conditional beliefs, and the levels $\alpha_{i}$ are sufficiently large, indicating a high level of risk tolerance.
 It is worth mentioning that the results in Proposition 3 can be explained from the perspective of Co-risk measures.
@@
 In the regulatory context, those $B_{i}$ could be stressing test events given by different supervising agencies.
-3.4 Absolutely continuous beliefs
-Next, we consider a different relation among the probability measures, absolute continuity. Due to the complexity of the problem, we here only consider the case $n=2$.
+Remark.
+The law-invariance of the inf-convolution for some risk functionals was discussed in Liu et al.
+(2020), showing that the law-invariance of the inf-convolution under homogeneous beliefs is guaranteed by some mild condition on the risk functionals if they are law-invariant.
+Proposition 4 shows that the inf-convolution of $\Lambda\mathrm{VaR}$ under homogeneous beliefs is law-invariant.
+However, if the beliefs are heterogeneous, the conclusion may not hold, which can be seen from Corollary 1.
+Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$ and $\alpha_{i}\in(0,1)$ satisfying $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}<1$.
+By Corollary 1, we have
+$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}}^{\mathbb{Q}}(X),$
+where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$.
+One can easily check that $X\overset{\mathbb{Q}_{i}}{=}Y$ for all $i=1,\dots,n$ does not imply $X\overset{\mathbb{Q}}{=}Y$.
+Hence, $\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)\neq\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(Y)$ may hold.
+3.4 General beliefs with $n=2$
+Next, we consider the case $n=2$. By Theorem A.17 of Föllmer and Schied (2016), for probability measures $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, we have
+$\mathbb{Q}_{2}(B)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B})+\mathbb{Q}_{2}(B\cap N)~\text{for all}~B\in\mathcal{F},$ (4)
+where $N\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(N)=0$ and $\eta=\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}\mathds{1}_{N^{c}}$.
+If $\mathbb{Q}_{2}(N)=0$, then $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ on $\Omega$; if $\mathbb{Q}_{2}(N)>0$, then $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ on $N^{c}$ but not on $N$.
+In this subsection, the relation of $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ is represented by (4).
 Theorem 4.
-Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ and denote $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ by $\eta$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
+For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
 $\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:g_{x}((1-F_{X}^{\mathbb{Q}_{1}}(x)-\Lambda_{1}(y))_{+})\leqslant\Lambda_{2}(x-y),~\text{for some}~y\in\mathbb{R}\right\},$
-where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t})=t$.
+where
+$g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)],$
+with $x_{t}$ satisfying $\mathbb{Q}_{1}(\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\})=t$.
 Moreover, if $x^{*}=\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)>$-$\infty$ and $g_{x^{*}}((1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+})\leqslant\Lambda_{2}(x^{*}-y^{*})$ for some $y^{*}\in\mathbb{R}$, then the optimal allocation is given by
 $X_{1}=\left(X-x^{*}\right)\mathds{1}_{\Omega\setminus A^{*}}+y^{*},~X_{2}=\left(X-x^{*}\right)\mathds{1}_{A^{*}}+x^{*}-y^{*},$ (5)
-where $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}$ with $z^{*}$ satisfying $\mathbb{Q}_{1}(X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*})=(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}$.
+where $A^{*}=\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}\cap N^{c}$ with $z^{*}$ satisfying $\mathbb{Q}_{1}(\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\})=(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}$.
+An immediate conclusion from Theorem 4 is the following result, showing that the inf-convolution is -$\infty$ if $\mathbb{Q}_{2}(N)$ is large enough.
+Corollary 2.
+Under the assumption of Theorem 4, if $\mathbb{Q}_{2}(N)\geqslant 1-\lambda_{2}^{-}$, then
+$\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)=-\infty.$
+Note that the idea used in the proof of Theorem 4 is not valid for the case $n\geqslant 3$. For the case $n\geqslant 3$, it may require a new method to simplify the main result obtained in Theorem 1.
 In Theorem 4, we do not obtain explicit formulas as the previous propositions.
@@
   1. (i)
-    If $X$ and $\eta$ are comonotonic, then $g_{x}(t)=\int_{F_{X}^{\mathbb{Q}_{1}}(x)}^{F_{X}^{\mathbb{Q}_{1}}(x)+t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$;
+    If $X$ and $\eta$ are comonotonic, then $g_{x}(t)=\int_{F_{X}^{\mathbb{Q}_{1}}(x)}^{F_{X}^{\mathbb{Q}_{1}}(x)+t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}\cap N^{c}$;
   2. (ii)
-    If $X$ and $\eta$ are countermonotonic, then $g_{x}(t)=\int_{0}^{t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\}$;
+    If $X$ and $\eta$ are countermonotonic, then $g_{x}(t)=\int_{0}^{t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\}\cap N^{c}$;
   3. (iii)
-    If $X$ and $\eta$ are independent under $\mathbb{Q}_{1}$, then $g_{x}(t)=(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{t/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ with the convention that $/0=0$ 0 and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant\frac{(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}}{1-F_{X}^{\mathbb{Q}_{1}}(x^{*})}\}$ with $U_{\eta}^{\mathbb{Q}_{1}}$ being independent of $X$ under $\mathbb{Q}_{1}$.
+    If $X$ and $\eta$ are independent under $\mathbb{Q}_{1}$, then $g_{x}(t)=(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{t/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ with the convention that $0/0=0$ and
+    $A^{*}=\{X>x^{*}\}\cap\left\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant\frac{(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}}{1-F_{X}^{\mathbb{Q}_{1}}(x^{*})}\right\}\cap N^{c}$
+    with $U_{\eta}^{\mathbb{Q}_{1}}$ being independent of $X$ under $\mathbb{Q}_{1}$.
 Note that the optimal set $A^{*}$ in (i) and (ii) have very different structures.
-To demonstrate this, suppose both $X$ and $\eta$ are continuous random variables under $\mathbb{Q}_{1}$.
+To demonstrate this, suppose both $X$ and $\eta$ are continuous random variables under $\mathbb{Q}_{1}$ and $N=\emptyset$.
 In (i), $A^{*}$ can be rewritten as $A^{*}=\{F_{X}^{\mathbb{Q}_{1}}(x^{*})<U_{X}^{\mathbb{Q}_{1}}<(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$.
@@
 Corollary 3.
-Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ and denote $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ by $\eta$. For $\alpha_{1},\alpha_{2}\in(0,1)$, we have
+For $\alpha_{1},\alpha_{2}\in(0,1)$, we have
 $\mathrm{VaR}_{\alpha_{1}}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{VaR}_{\alpha_{2}}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:g_{x}((1-F_{X}^{\mathbb{Q}_{1}}(x)-\alpha_{1})_{+})\leqslant\alpha_{2}\right\},$
-where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t})=t$.
+where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\})=t$.
 Under the assumption of Corollary 3, we have the following results.
@@
     $\mathrm{VaR}_{\alpha_{1}}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{VaR}_{\alpha_{2}}^{\mathbb{Q}_{2}}(X)=\inf\left\{x\in\mathbb{R}:(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{(1-F_{X}^{\mathbb{Q}_{1}}(x)-\alpha_{1})_{+}/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s\leqslant\alpha_{2}\right\}$
-    with the convention that $/0=0$ 0.
-4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogenous beliefs
-Throughout this section, we set $\mathcal{X}=L^{\infty}$.
-We investigate the risk sharing problem between two agents, considering the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and a monotone risk measure $\rho^{\mathbb{Q}_{2}}$, where $\mathbb{Q}_{2}$ may be different from $\mathbb{Q}_{1}$.
+    with the convention that $0/0=0$.
+4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogeneous beliefs
+In this section, we investigate the risk sharing problem between two agents, considering the inf-convolution of $\Lambda\mathrm{VaR}$ and a monotone risk measure $\rho$ under the probability measures $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, where $\mathbb{Q}_{2}$ may be different from $\mathbb{Q}_{1}$.
 Note that $\rho$ here is very general including many commonly used risk measures as special cases.
+In our results below, we consider both the cases $\mathcal{X}=L^{\infty}$ and unbounded $\mathcal{X}$, each with different assumptions on $\rho$.
 Theorem 5.
-Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\rho$ is monotone. Then we have
+Suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ and $\rho$ is monotone.
+If either (i) $\mathcal{X}=L^{\infty}$ and $\rho$ is law-invariant under $\mathbb{Q}_{2}$ or (ii) $\mathcal{X}$ is unbounded and $\rho$ is an $\varepsilon$-tail risk measure under $\mathbb{Q}_{2}$ for some $\varepsilon\in(0,1)$, then we have
 $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$ (6)
@@
 The result of Theorem 5 is still new even if $\Lambda\mathrm{VaR}$ is reduced to $\mathrm{VaR}$.
-Let us now consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$.
+Let us now consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$, where $\mathbb{P}$ is atomless.
 Under this setup, the contingent set $B$ in (6) is given explicitly.
@@
 Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~i=1,2$.
-Moreover, suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\rho$ is monotone.
-Then we have
+Moreover, suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ and $\rho$ is monotone.
+If either (i) $\mathcal{X}=L^{\infty}$ and $\rho$ is law-invariant under $\mathbb{Q}_{2}$ or (ii) $\mathcal{X}$ is unbounded and $\rho$ is an $\varepsilon$-tail risk measure under $\mathbb{Q}_{2}$ for some $\varepsilon\in(0,1)$, then we have
 $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\},$ (8)
@@
 Let $\mathcal{G}=\{g:[0,1]\to[0,1]|~g~\text{is increasing and left-continuous satisfying}~g(0)=0~\text{and}~g(1)=1\}$.
-For $g\in\mathcal{H}$, the distortion risk measure $\rho_{g}^{\mathbb{Q}}$ is defined as
+For $g\in\mathcal{H}$, the distortion risk measure $\rho_{g}$ under $\mathbb{Q}$ is defined as
 $\rho_{g}^{\mathbb{Q}}(X)$ $=\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
@@
 (2022).
-Furthermore, for an increasing function $u$, the expected utility is defined as $\mathbb{E}_{u}^{\mathbb{Q}}:X\mapsto\mathbb{E}^{\mathbb{Q}}(u(X))$.
 In the following corollary, let $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{1}\cap B_{2})>0$, indicating that beliefs are not totally incompatible, and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|B_{1}\cap B_{2})$, which in a sense represents the consensus view of the two agents.
-Applying Proposition 5, we obtain the inf-convolution of $\mathrm{\Lambda\mathrm{VaR}}$ and $\rho_{g}/\mathbb{E}_{u}/\Lambda_{1}\mathrm{VaR}^{+}$ under heterogeneous beliefs, respectively.
+Applying Proposition 5, we obtain the inf-convolution of $\mathrm{\Lambda\mathrm{VaR}}$ and $\rho_{g}/\Lambda_{1}\mathrm{VaR}^{+}$ under heterogeneous beliefs, respectively.
 In order to simplify the expression, we suppose $\lambda^{+}$ is attainable for $\Lambda$, i.e., there is $x\in\mathbb{R}$ such that $\Lambda(x)=\lambda^{+}$.
 Corollary 5.
-Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\lambda^{+}$ is attainable.
+Suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ and $\lambda^{+}$ is attainable.
 Then the following conclusion holds with $\beta=\frac{(1-\lambda^{+})\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$.
   1. (i)
-    For $g\in\mathcal{G}$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$ then
+    For $g\in\mathcal{G}$ satisfying $g(\varepsilon)=1$ for some $\varepsilon\in(0,1)$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=$-$\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$ then
     $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=\left\{\begin{array}[]{cc}\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t),&g(\beta)=1\\ -\infty,&g(\beta)<1\end{array}\right$.;
   2. (ii)
-    For an increasing function $u$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
-    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\left(x+(1-\beta_{x})u(-\infty)+\int_{0}^{\beta_{x}}u\left(\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)-x\right)\mathrm{d}t\right),$
-    where $u(-\infty)=\lim_{y\to-\infty}u(y)$ and $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$;
-  3. (iii)
-    For $\Lambda_{1}\in\mathcal{H}$ with $10<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<$ 1, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
+    For $\Lambda_{1}\in\mathcal{H}$ with $0<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<1$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=$-$\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$, then
     $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\overline{\Lambda}^{x}\mathrm{VaR}^{+,\mathbb{Q}}(X),$
@@
 For instance, (i) of Corollary 5 exactly extends Theorem 5.3 of Wang and Wei (2018) ($\rho$ is a distortion risk measure which is law-invariant under $\mathbb{Q}_{1}$) in these two directions.
-It is worth mentioning that (iii) of Corollary 4 is a new result even for the case $\mathbb{Q}_{1}=\mathbb{Q}_{2}$.
-Let us now consider the case $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$.
+It is worth mentioning that (ii) of Corollary 5 is a new result even for the case $\mathbb{Q}_{1}=\mathbb{Q}_{2}$.
+Next, suppose $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ has the relation (4).
+For an increasing function $u$, the expected utility is defined as $\mathbb{E}_{u}^{\mathbb{Q}}:X\mapsto\mathbb{E}^{\mathbb{Q}}(u(X))$.
 Instead of finding a simplified version of Theorem 5 for this case, we apply Theorem 5 directly to some specific $\rho$.
@@
 Corollary 6.
-Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ with $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ denoted by $\eta$.
-For $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1, we have the following conclusions.
+For $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$, we have the following conclusions.
   1. (i)
@@
   2. (ii)
-    For an increasing function $u$, we have
-    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\left(x+u(-\infty)\Lambda(x)\mathrm{ES}_{\Lambda(x)}^{\mathbb{Q}_{1}}(\eta)+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})\right),$
+    For $\mathcal{X}=L^{\infty}$ and an increasing function $u$, we have
+    $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$
+    $=\inf_{x\in\mathbb{R}}\left(x+u(-\infty)(\mathbb{Q}_{2}(N)+\Lambda(x)\mathrm{ES}_{\Lambda(x)}^{\mathbb{Q}_{1}}(\eta))+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})\right),$
     where $u(-\infty)=\lim_{y\to-\infty}u(y)$.
@@
 (2020) for $n=2$, where the semi-explicit formula is offered.
-For the case $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$, it might be difficult to find more explicit expression for the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and $\rho_{g}^{\mathbb{Q}_{2}}$/$\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$.
+Using (4), it might be difficult to find more explicit expression for the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and $\rho_{g}^{\mathbb{Q}_{2}}$$\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$.
 5 Inf-convolution of $\Lambda\mathrm{VaR}^{+}$ and a monotone risk measure
@@
 (2018).
-In this section, we focus on the inf-convolution of $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}$ and a monotone risk measure $\rho^{\mathbb{Q}_{2}}$, which includes the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ as a special case.
-Let $\mathcal{X}=L^{\infty}$.
+In this section, we focus on the inf-convolution of $\Lambda\mathrm{VaR}^{+}$ and a monotone risk measure $\rho$ under $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, which includes the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ as a special case.
 For any $\Lambda\in\mathcal{H}$, we denote $\Lambda_{x}(z)=\inf_{x\leqslant t\leqslant z}\Lambda(t),~z\geqslant x$.
@@
 The following result is valid for general $\Lambda\in\mathcal{H}$.
+We say $\rho$ is continuous from above on $\mathcal{X}$ if $X_{n}\downarrow X$ implies $\lim_{n\to\infty}\rho(X_{n})=\rho(X)$ for $X_{n},X\in\mathcal{X}$.
 Theorem 6.
-Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1, and $\rho$ is monotone and law-invariant under $\mathbb{Q}_{2}$. Then we have
+Suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$, and $\rho$ is monotone and law-invariant under $\mathbb{Q}_{2}$.
+If either (i) $\mathcal{X}=L^{\infty}$ or (ii) $\mathcal{X}$ is unbounded and $\rho$ is continuous from above, then we have
 $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\},$ (10)
@@
 $\Lambda_{x,y}(z)=\left\{\begin{array}[]{cc}0,&z<x\\ 1-\Lambda_{x}(z),&x\leqslant z<y\\ 1,&z\geqslant y\end{array}\right.$.
-Moreover, the optimal allocation of the inf-convolution exists if and only if the minimizer of (9) exists. If $(x^{*},y^{*},U^{*})$ is the minimizer, then one optimal allocation is given by
+Moreover, for case (i), the optimal allocation of the inf-convolution exists if and only if the minimizer of (10) exists.
+For both cases (i) and (ii), if $(x^{*},y^{*},U^{*})$ is the minimizer, then one optimal allocation is given by
 $X_{1}^{*}=\Lambda_{x^{*},y^{*}}^{-1}(U^{*}),~X_{2}^{*}=X-\Lambda_{x^{*},y^{*}}^{-1}(U^{*}).$ (11)
-In the case of $\rho=\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$ for $\Lambda_{1}\in\mathcal{H}$ with $10<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<$ 1, it follows from Theorem 6 that the optimal risk allocation for risk sharing between $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}$ and $\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}$ has the form of (10), which is neither comonotonic nor pairwise countermonotonic.
-This means that the optimal risk sharing for multiple $\Lambda\mathrm{VaR}^{+,\mathbb{Q}}$ is a very different problem from the cases with convex risk measures or quantile-based risk measures in the literature.
+In the case of $\rho=\Lambda_{1}\mathrm{VaR}^{+}$ for $\Lambda_{1}\in\mathcal{H}$ with $0<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<1$, it follows from Theorem 6 that the optimal risk allocation for risk sharing between $\Lambda\mathrm{VaR}^{+}$ and $\Lambda_{1}\mathrm{VaR}^{+}$ under $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ has the form of (11), which is neither comonotonic nor pairwise countermonotonic.
+This means that the optimal risk sharing for multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is qualitatively different from the cases with convex risk measures or quantile-based risk measures in the literature.
 The expression of the inf-convolution in (10) is an optimization problem involving three parameters $(x,y,U)$, where $U\sim U[0,1]$ under $\mathbb{Q}_{1}$.
@@
 Moreover, the expression in (10) and also the optimal allocation in (11) show that shape of $\Lambda$ function plays an important role in both the expression of the inf-convolution and the form of the optimal allocations, which is very different from the results of $\Lambda\mathrm{VaR}$ displayed in Sections 3 and 4.
+For case (ii) of Theorem 6, we suppose that $\rho$ is continuous from above.
+In fact, many risk functionals satisfy this assumption, such as distortion risk measures $\rho_{g}$ with continuous $g$ (including $\mathrm{ES}_{\alpha}$ as a special case), expected utility $\mathbb{E}_{u}$ for continuous and increasing $u$, and the right quantiles.
+Later, we will define an alternative Lambda-$\mathrm{VaR}$ measure that also satisfies this assumption.
 We are also aware that the conclusion in Theorem 6 is complicated in the sense that it involves $\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x,y}^{-1}(U))$, which is a very difficult problem even if $\mathbb{Q}_{1}=\mathbb{Q}_{2}$.
@@
 (2024), Han and Liu (2025) and the references therein for the discussion on the robust risk aggregation.
+We denote $X\leqslant_{\mathrm{i}cx}Y$ if $\mathbb{E}(f(X))\leqslant\mathbb{E}(f(Y))$ for all increasing and convex function $f$.
+A mapping $\rho:\mathcal{X}\to\mathbb{R}$ is said to be SSD-consistent if $X\leqslant_{\mathrm{i}cx}Y$ implies $\rho(X)\leqslant\rho(Y)$.
+Note that here $\mathcal{X}\subset L^{1}$.
+Corollary 7.
+Suppose $\mathcal{X}\subset L^{1}$. If $\mathbb{Q}_{1}=\mathbb{Q}_{2}=\mathbb{Q}$ and $\rho$ is SSD-consistent risk measure, then under the assumption of Theorem 6, we have
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\left\{x+\rho^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U_{X}^{\mathbb{Q}})\right)\right\}.$
+Note that the above corollary coincides with Theorem 5 of Liu (2025).
+We next see an example that $\rho$ is not SSD-consistent.
+For $\Lambda:\mathbb{R}\to[0,1]$ and an increasing $f:\mathbb{R}\to[0,1]$, define
+$q_{\Lambda}^{+}(f)=\inf\{x\in\mathbb{R}:f(x)>1-\Lambda(x)\},$
+where $\inf\emptyset=\infty$.
+Note that if $\Lambda$ is a constant and $f$ is a distribution function, then $q_{\Lambda}^{+}(f)$ is the right quantile.
+We refer to Bellini and Peri (2022) and Han and Liu (2025) for the properties and applications of $q_{\Lambda}^{+}$.
+Define an alternative version of $\Lambda\mathrm{VaR}$ by $\overline{\Lambda\mathrm{VaR}}(X)=q_{\Lambda}^{+}(F_{X})$ for $X\in\mathcal{X}$.
+One can easily show that $\overline{\Lambda\mathrm{VaR}}$ is continuous from above if $\Lambda\in\mathcal{H}$.
+Then applying Theorem 6, we have the following result.
+Corollary 8.
+Suppose $\mathbb{Q}_{1}=\mathbb{Q}_{2}=\mathbb{Q}$, $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ is strictly decreasing and $\Lambda_{1}\in\mathcal{H}$ with $0<\lambda_{1}^{-}\leqslant\lambda_{1}^{+}<1$.
+Then we have, for $X\in\mathcal{X}$ with continuous $F_{X}^{\mathbb{Q},-1}$ over $(0,1)$,
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}\mathop{\square}\displaylimits\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\left\{x+q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})\right\},$
+where $l_{x,y}(t)=\sup_{s\in(0,t]}\{F_{X}^{\mathbb{Q},-1}(s)-\Lambda_{x,y}^{-1}(1-t+s)\},~t\in(0,1)$ and $l_{x,y}^{-1}(z)=\inf\{w\in(0,1):l_{x,y}(w)\geqslant z\},~z\in\mathbb{R}$ with $\inf\emptyset=1$.
+Corollary 8 requires that $F_{X}^{\mathbb{Q},-1}$ is continuous over $(0,1)$, which is a technical assumption for the computation of $\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U)\right)$.
+Our computation may not be correct if this assumption is violated.
 In order to simplify (10), we consider a special case of $\Lambda$ functions. This makes it easier to find more explicit expressions of the inf-convolution for concrete examples of $\rho$.
 Proposition 6.
-For $x_{1}\in\mathbb{R}$ and $10<\lambda_{1}<\lambda_{2}<$ 1, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$.
+For $x_{1}\in\mathbb{R}$ and $0<\lambda_{1}<\lambda_{2}<1$, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$.
 Under the assumption of Theorem 6, we have
 $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-x\mathds{1}_{B_{1}}-x\vee x_{1}\mathds{1}_{B_{2}}-y\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)\right\},$ (12)
-where $\mathcal{B}_{\lambda_{1},\lambda_{2}}=\{(B_{1},B_{2}):\mathbb{Q}_{1}(B_{1})=\lambda_{1},\mathbb{Q}_{2}(B_{2})=\lambda_{2}-\lambda_{1},B_{1}\cap B_{2}=\emptyset\}$.
+where $\mathcal{B}_{\lambda_{1},\lambda_{2}}=\{(B_{1},B_{2}):\mathbb{Q}_{1}(B_{1})=\lambda_{1},\mathbb{Q}_{1}(B_{2})=\lambda_{2}-\lambda_{1},B_{1}\cap B_{2}=\emptyset\}$.
 Note that the optimization problem in (12) is much easier than that in (10) because it only has parameters $(x,y,B_{1},B_{2})$.
+To illustrate our Proposition 6, we next see the example of $\rho^{\mathbb{Q}_{2}}=\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}$.
+In the following corollary, we suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|D_{i})$ for some $D_{i}\in\mathcal{F},~i=1,2$, and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|D_{1}\cap D_{2})$ with $\mathbb{P}(D_{1}\cap D_{2})>0$.
+Corollary 9.
+For $x_{1}\in\mathbb{R}$ and $0<\lambda_{1}<\lambda_{2}<1$, let $\Lambda(z)=(1-\lambda_{1})\mathds{1}_{\{z<x_{1}\}}+(1-\lambda_{2})\mathds{1}_{\{z\geqslant x_{1}\}}$. Then we have
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\left\{\begin{array}[]{cc}\frac{1}{\alpha}\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}\int_{\alpha_{0}}^{\alpha_{1}}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t,&\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})\geqslant\alpha\mathbb{P}(D_{2})\\ -\infty,&\text{otherwise}\end{array}\right$.,
+where $\alpha_{0}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})}{\mathbb{P}(D_{1}\cap D_{2})}$ and $\alpha_{1}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})-\alpha\mathbb{P}(D_{2})}{\mathbb{P}(D_{1}\cap D_{2})}$.
 6 Conclusion
-In this paper, we consider the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under heterogenous beliefs.
+In this paper, we consider the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under heterogeneous beliefs.
 We obtain the general expression for the inf-convolution and the corresponding optimal allocations.
-To study impact of the relation of the beliefs on the inf-convolution, we discuss three different cases: i) homogeneous beliefs; ii) conditional beliefs; iii) absolute continuous beliefs.
+To study impact of the relation of the beliefs on the inf-convolution, we discuss three different cases: i) homogeneous beliefs; ii) conditional beliefs; iii) general beliefs with two agents.
 For all these cases, we obtain more explicit expressions of the inf-convolution.
 In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high.
-The inf-convolution of one ΛVaR/$\Lambda\mathrm{VaR}/\Lambda\mathrm{VaR}^{+}$ and a general risk measures with belief heterogeneity are also discussed.
-There are still some unsolved problems such as the inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs, which deserves further study.
-Acknowledgements: YW acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2023-04674).
-7 Declarations
-Conflict of interest: All authors declare that they have no conflict of interest.
-Appendix A Proof of Section 2
+The inf-convolution of one $\Lambda\mathrm{VaR}/\Lambda\mathrm{VaR}^{+}$ and a general risk measure with belief heterogeneity are also discussed.
+There are still some unsolved problems, such as what the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under general heterogeneous beliefs is, and what the inf-convolution of multiple $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs is, which deserves further study.
+Acknowledgments
+The authors thank the editor, an associate editor, and two anonymous referees for their helpful comments.
+Competing Interests
+The authors declare no competing interests.
+Appendix A Proof of all results in Sections 2-5
 Proof of Proposition 1.
@@
 as $c\to$-$\infty$. Hence, $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)=-\infty.$ This completes the proof. ∎
-Appendix B Proof of Section 3
-In this section, we give all of the proofs of the results in Section 3.
 Proof of Theorem 1.
@@
 For $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, let $x=\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$ and $y_{i}=\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$.
-Note that $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 implies $y_{i}\in\mathbb{R}$ and $x\in\mathbb{R}$.
+Note that $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$ implies $y_{i}\in\mathbb{R}$ and $x\in\mathbb{R}$.
 By the definition of $\Lambda_{i}\mathrm{VaR}$ and the right continuity of $\Lambda_{i}$, we have $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$ with $C_{i}=\{X_{i}>y_{i}\}$.
@@
 For $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, let $A_{i}\subset D_{i}$. Then we have
-$\{X>x,A_{i}\}\subset\{X>x,D_{i}\}=\{X>x,C_{i}\}\subset C_{i}.$
-Using the fact $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$, we have $\mathbb{Q}_{i}(X>x,A_{i})\leqslant\Lambda(y_{i})$.
+$\{X>x\}\cap A_{i}\subset\{X>x\}\cap D_{i}=\{X>x\}\cap C_{i}\subset C_{i}.$
+Using the fact $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$, we have $\mathbb{Q}_{i}(\{X>x\}\cap A_{i})\leqslant\Lambda(y_{i})$.
 This implies $\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})=x\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
@@
 Next we show the inverse inequality.
-Suppose there exist $y_{i}\in\mathbb{R},(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$.
+Suppose there exist $y_{i}\in\mathbb{R},(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})$.
 Let $X_{i}=(X-x)\mathds{1}_{A_{i}}+y_{i},~i=1,\dots,n$, where $x=\sum_{i=1}^{n}y_{i}$.
@@
 where $\Lambda_{i}^{y_{i}}(z)=\Lambda_{i}(z+y_{i})$ for $z\in\mathbb{R}$.
-Note that $\mathbb{Q}_{i}((X-x)\mathds{1}_{A_{i}}>0)=\mathbb{Q}_{i}(X>x,A_{i})\leqslant\Lambda_{i}(y_{i})=\Lambda_{i}^{y_{i}}(0)$.
+Note that $\mathbb{Q}_{i}((X-x)\mathds{1}_{A_{i}}>0)=\mathbb{Q}_{i}(\{X>x\}\cap A_{i})\leqslant\Lambda_{i}(y_{i})=\Lambda_{i}^{y_{i}}(0)$.
 By definition, we have $\Lambda_{i}^{y_{i}}\mathrm{VaR}^{\mathbb{Q}_{i}}((X-x)\mathds{1}_{A_{i}})\leqslant 0$.
@@
 Note that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\leqslant 1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ implies $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{-}$ for all $i=1,\dots,n$.
-Hence, for any $(y_{1},\dots,y_{n})$, it follows that $\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$.
+Hence, for any $(y_{1},\dots,y_{n})$, it follows that $\mathbb{Q}_{i}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i}\right)\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{-}\leqslant\Lambda_{i}(y_{i})$.
 Consequently, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
 Case (ii).
-Let $x^{*}=\bigwedge_{i=1}^{n}\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}X$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ represents the $\mathrm{ess\mbox{-}inf}$ under $\mathbb{Q}_{i}$.
-For any $(y_{1},\dots,y_{n})$ with $\sum_{i=1}^{n}y_{i}<x^{*}$ and $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, there exists $i\in\{1,\dots,n\}$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}^{n}y_{i},A_{i})=\mathbb{Q}_{i}(A_{i})>\lambda_{i}^{+}\geqslant\Lambda_{i}(y_{i})$.
-Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)\geqslant x^{*}>-\infty$.
-∎
+We assume by contradiction that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
+By definition, there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ with $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$ and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(\{X>\sum_{j=1}^{n}y_{j}^{(m)}\}\cap A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$ for $i=1,\dots,n$.
+Moreover, using the condition $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$, there exists $\varepsilon>0$ such that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1+\varepsilon$ for any $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$.
+Hence, there exists $i_{m}\in\{1,2,\dots,n\}$ such that $\mathbb{Q}_{i_{m}}(A_{i_{m}}^{(m)})>(1+\varepsilon)\lambda_{i_{m}}^{+}$.
+For $\{i_{m},m\geqslant 1\}$, we can find a constant subsequence.
+Without loss of generality, we assume $i_{m}=1$ for all $m\geqslant 1$.
+This means $\mathbb{Q}_{1}(A_{1}^{(m)})>(1+\varepsilon)\lambda_{1}^{+}$.
+This together with the fact that $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$ implies for $m$ large enough,
+$\mathbb{Q}_{1}\left(\left\{X>\sum_{j=1}^{n}y_{j}^{(m)}\right\}\cap A_{1}^{(m)}\right)$ $=\mathbb{Q}_{1}(A_{1}^{(m)})-\mathbb{Q}_{1}\left(\left\{X\leqslant\sum_{j=1}^{n}y_{j}^{(m)}\right\}\cap A_{1}^{(m)}\right)$
+$>(1+\varepsilon)\lambda_{1}^{+}-\mathbb{Q}_{1}\left(X\leqslant\sum_{j=1}^{n}y_{j}^{(m)}\right)$
+$\geqslant(1+\varepsilon/2)\lambda_{1}^{+},$
+which contradicts $\mathbb{Q}_{1}(\{X>\sum_{j=1}^{n}y_{j}^{(m)}\}\cap A_{1}^{(m)})\leqslant\Lambda_{1}(y_{1}^{(m)})\leqslant\lambda_{1}^{+}$.
+Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
 Proof of Proposition 3.
@@
 Hence, there exists $y_{i}^{0}$ such that $\Lambda(y_{i})\geqslant\mathbb{Q}_{i}(A_{i})$ for all $y_{i}<y_{i}^{0}$.
-It follows that $\mathbb{Q}_{i}(X>\sum_{i=1}^{n}y_{i},A_{i})\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\Lambda_{i}(y_{i})$ if $y_{i}\leqslant y_{i}^{0}$ for all $i=1,\dots,n$.
+It follows that $\mathbb{Q}_{i}(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i})\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\Lambda_{i}(y_{i})$ if $y_{i}\leqslant y_{i}^{0}$ for all $i=1,\dots,n$.
 This implies $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
@@
 Note that there exists $y_{j}^{0}\in\mathbb{R}$ such that $\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j}^{0})$ for all $j\neq i$.
-Moreover, for any $y_{i}\in\mathbb{R}$, it follows that $\mathbb{Q}_{j}(X>y_{i}+\sum_{j\neq i}y_{j}^{0},A_{j})\leqslant\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j})$ for $j\neq i$, and $\mathbb{Q}_{i}(X>y_{i}+\sum_{j\neq i}y_{j}^{0},A_{i})\leqslant\Lambda_{i}(y_{i})$.
+Moreover, for any $y_{i}\in\mathbb{R}$, it follows that $\mathbb{Q}_{j}(\{X>y_{i}+\sum_{k\neq i}y_{k}^{0}\}\cap A_{j})\leqslant\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j}^{0})$ for $j\neq i$, and $\mathbb{Q}_{i}(\{X>y_{i}+\sum_{k\neq i}y_{k}^{0}\}\cap A_{i})\leqslant\Lambda_{i}(y_{i})$.
 By the arbitrary of $y_{i}$, we have $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
 We next show the last statement.
-We denote $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)=1+\varepsilon$ for some $\varepsilon>0$.
+Let $1+\varepsilon=\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)$ for some $\varepsilon>0$.
 Suppose by contradiction that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
-Then there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ satisfying $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=-\infty$, and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}y_{i}^{(m)},A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$.
+Then there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ satisfying $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$, and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(\{X>\sum_{j=1}^{n}y_{j}^{(m)}\}\cap A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$.
 Without loss of generality, we can assume $\lim_{m\to\infty}y_{i}^{(m)}=y_{i}$.
@@
 By the fact that $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$, we have $E_{1}\neq\emptyset$.
-Moreover, there exists $m_{0}>0$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}y_{i}^{(m)})=1$ for all $m\geqslant m_{0}$.
-Hence, we have $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$ for all $m\geqslant m_{0}$.
+Moreover, for any $0<\eta<\varepsilon$, there exists $m_{0}>0$ such that $\mathbb{Q}_{i}(X\leqslant\sum_{j=1}^{n}y_{j}^{(m)})<\lambda_{i}^{-}\eta/2$ for all $m\geqslant m_{0}$.
+Hence, we have $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant(1+\eta/2)\Lambda_{i}(y_{i}^{(m)})$ for all $m\geqslant m_{0}$.
 Note that $\lim_{m\to\infty}\Lambda_{i}(y_{i}^{(m)})=\lambda_{i}^{-}$ for all $i\in E_{1}$.
 Hence, for any $0<\eta<\varepsilon$, there exists $m_{1}\geqslant m_{0}$ such that $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{-}(1+\eta)$ for all $m\geqslant m_{1}$ and $i\in E_{1}$.
-Moreover, $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{+}$ for all $m\geqslant m_{1}$ and $i\in E_{2}$.
-Combination of the those conclusion leads to $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i}^{(m)})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)\leqslant 1+\eta<1+\varepsilon$ for $m>m_{1}$, which is a contradiction of the assumption.
+Moreover, $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{+}(1+\eta)$ for all $m\geqslant m_{1}$ and $i\in E_{2}$.
+Combination of the those conclusion leads to $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i}^{(m)})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j}^{(m)})}{\lambda_{j}^{+}}\right)\leqslant 1+\eta<1+\varepsilon$ for $m>m_{1}$, which is a contradiction of the assumption.
 Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
@@
 Proof of Theorem 2. By Theorem 1, we have
-$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
-Note that $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ if and only if $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})$.
+$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
+Note that $\mathbb{P}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ if and only if $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})$.
 Hence, we have
@@
 For the optimal allocation, note that the right continuity of $\Lambda^{*}$ at $x^{*}$ implies $\mathbb{P}\left(X>x^{*}\right)\leqslant\Lambda^{*}(x^{*})$.
-Moreover, there exist $(y_{1}^{*},\dots,y_{n}^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ such that $1\bigwedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $\mathbb{P}\left(X>x^{*},A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
+Moreover, there exist $(y_{1}^{*},\dots,y_{n}^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ such that $1\bigwedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $\mathbb{P}\left(\{X>x^{*}\}\cap A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
 Hence, $(y_{1}^{*},\dots,y_{n}^{*},A_{1}^{*},\dots,A_{n}^{*})$ is the minimizer of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
@@
 $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$
-$=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
+$=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
 Let $\mathcal{N}_{n}=\{\{i_{1},\dots,i_{m}\}\subset\{1,\dots,n\}:~m=1,2,\dots,n-1\}$.
@@
 Hence, we have
-$\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}$
-$\leqslant\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
+$\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}$
+$\leqslant\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap\left(\cap_{j=1}^{n}B_{j}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
 Note that the inverse inequality is trivial. Hence, we have
 $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$
-$=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
+$=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap\left(\cap_{j=1}^{n}B_{j}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
 Clearly, $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=$-$\infty$ if $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)=0$.
 Next, we consider the case $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)>0$.
-Note that $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ and some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ is equivalent to $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\cap_{i=1}^{n}B_{i}\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$.
+Note that $\mathbb{P}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap\left(\cap_{j=1}^{n}B_{j}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ and some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ is equivalent to $\mathbb{P}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap(\cap_{j=1}^{n}B_{j})\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$.
 Using this conclusion and the attainability of $\Lambda^{\diamond}$, we have
-$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$ $=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{P}\left(X>x,\cap_{i=1}^{n}B_{i}\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})\right\}$
-$=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{Q}\left(X>x\right)\leqslant\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\}$
+$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$ $=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{P}\left(\{X>x\}\cap(\cap_{j=1}^{n}B_{j})\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})\right\}$
+$=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{Q}\left(X>x\right)\leqslant\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}\right\}$
 $=\inf\left\{x:\mathbb{Q}\left(X>x\right)\leqslant\Lambda^{\diamond}(x)\right\}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X).$
@@
 $\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)$
-$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{Q}_{2}\left(X>x,A\right)\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},A\in\mathcal{F}\right\}$
-$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A\}})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},A\in\mathcal{F}\right\}.$
-Let $f(t)=\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant t),~{}t\in[0,1]$.
+$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{Q}_{2}\left(\{X>x\}\cap A\right)\leqslant\Lambda_{2}(x-y),~\text{for some}~y\in\mathbb{R},A\in\mathcal{F}\right\}$
+$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A})+\mathbb{Q}_{2}(\{X>x\}\cap A\cap N)\leqslant\Lambda_{2}(x-y),\right$.
+$~~~~~~~~~~~~\left.\text{for some}~y\in\mathbb{R},A\in\mathcal{F}\right\}.$
+Let $A^{\prime}=A\cap N^{c}$.
+Using the fact $\mathbb{Q}_{1}(N)=0$, we have $\mathbb{Q}_{1}\left(\{X>x\}\cap(A^{\prime})^{c}\right)=\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)$, $\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A^{\prime}})=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A})$ and $\mathbb{Q}_{2}(\{X>x\}\cap A^{\prime}\cap N)=0$.
+Hence, the inf-convolution can be simplified as
+$\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)$
+$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A})\leqslant\Lambda_{2}(x-y),~\text{for some}~y\in\mathbb{R},A\in\mathcal{F}\right\}.$
+Let $f(t)=\mathbb{Q}_{1}(\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant t\}),~t\in[0,1]$.
 Clearly, $f$ is an increasing and continuous function.
@@
 Define $A_{x,t}=\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant f^{-1}(t)\}$ for $t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$.
-Let $t_{0}=\mathbb{Q}_{1}(X>x,A)$.
-Then it follows that $\mathbb{Q}_{1}(X>x,A_{x,t_{0}})=\mathbb{Q}_{1}(X>x,A)$ and
-$\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A_{x,t_{0}}\}})=\mathbb{E}^{\mathbb{Q}_{1}}(F_{\eta}^{\mathbb{Q}_{1},-1}(U_{\eta}^{\mathbb{Q}_{1}})\mathds{1}_{\{X>x,A_{x,t_{0}}\}})\leqslant\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A\}}).$
+Let $t_{0}=\mathbb{Q}_{1}(\{X>x\}\cap A)$.
+Then it follows that $\mathbb{Q}_{1}(\{X>x\}\cap A_{x,t_{0}})=\mathbb{Q}_{1}(\{X>x\}\cap A)$ and
+$\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A_{x,t_{0}}})=\mathbb{E}^{\mathbb{Q}_{1}}(F_{\eta}^{\mathbb{Q}_{1},-1}(U_{\eta}^{\mathbb{Q}_{1}})\mathds{1}_{\{X>x\}\cap A_{x,t_{0}}})\leqslant\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A}).$
 Consequently, we have
 $\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)$
-$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A_{x,t}^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A_{x,t}\}})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
+$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A_{x,t}^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A_{x,t}})\leqslant\Lambda_{2}(x-y),\right$.
+$~~~~~~~~~~~~\left.\text{for some}~y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
 $=\inf\left\{x\in\mathbb{R}:1-F_{X}^{\mathbb{Q}_{1}}(x)-t\leqslant\Lambda_{1}(y),g_{x}(t)\leqslant\Lambda_{2}(x-y),~\text{for some}~y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
@@
 Next, we show $(X_{1},X_{2})$ given by (5) is the optimal allocation. Note that
-$\mathbb{Q}_{1}(X_{1}>y^{*})$ $=\mathbb{Q}_{1}(X>x^{*},\Omega\setminus A^{*})=\mathbb{Q}_{1}(X>x^{*})-\mathbb{Q}_{1}(X>x^{*},A^{*})$
+$\mathbb{Q}_{1}(X_{1}>y^{*})$ $=\mathbb{Q}_{1}(\{X>x^{*}\}\cap(\Omega\setminus A^{*}))=\mathbb{Q}_{1}(X>x^{*})-\mathbb{Q}_{1}(\{X>x^{*}\}\cap A^{*})$
 $=1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\leqslant\Lambda_{1}(y^{*}),$
-$\mathbb{Q}_{2}(X_{2}>x^{*}-y^{*})$ $=\mathbb{Q}_{2}(X>x^{*},A^{*})=\mathbb{E}^{\mathbb{Q}_{1}}\left(\eta\mathds{1}_{\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}}\right)$
+$\mathbb{Q}_{2}(X_{2}>x^{*}-y^{*})$ $=\mathbb{Q}_{2}(\{X>x^{*}\}\cap A^{*})=\mathbb{E}^{\mathbb{Q}_{1}}\left(\eta\mathds{1}_{\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}}\right)$
 $=g_{x^{*}}((1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+})\leqslant\Lambda_{2}(x^{*}-y^{*}).$
@@
 ∎
-Appendix C Proof of Section 4
-In this section, we provide all the proofs of results in Section 4.
+Proof of Corollary 2.
+Note that $\mathbb{Q}_{2}(N)\geqslant 1-\lambda_{2}^{-}$ implies $g_{x}(t)\leqslant 1-\mathbb{Q}_{2}(N)\leqslant\lambda_{2}^{-}\leqslant\Lambda_{2}(x-y)$ for all $x,y\in\mathbb{R}$.
+Hence, the inf-convolution is -$\infty$.
+∎
 Proof of Theorem 5. For any $x,y$ and $B$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, let $X_{1}=x\mathds{1}_{B}+(X-y)\mathds{1}_{B^{c}}$. Direct calculation gives
@@
 under $\mathbb{Q}_{1}$.
-Define $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ and $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{i}}$ represent the $\mathrm{ess\mbox{-}inf}$ and $\mathrm{ess\mbox{-}sup}$ under $\mathbb{Q}_{i}$, $i=1,2$.
+Define $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>m_{0}$.
 Note that $U_{X_{1}}^{\mathbb{Q}_{1}}$ can be chosen such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$.
 By the definition of $\Lambda\mathrm{VaR}$ and the right continuity of $\Lambda$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x_{1})\geqslant 1-\Lambda(x_{1})$, which implies $x_{1}\geqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(1-\Lambda(x_{1}))$.
-Hence we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(Y_{1})=x_{1}$, and $X_{1}\leqslant Y_{1}$ for all $w\in\Omega$.
-By monotonicity of $\rho$, we have
+Hence, $X_{1}\leqslant x_{1}$ for all $w\in\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}$.
+For the case $\mathcal{X}=L^{\infty}$, let $m_{0}=(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ and $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{i}}$ represent the $\mathrm{ess\mbox{-}inf}$ and $\mathrm{ess\mbox{-}sup}$ under $\mathbb{Q}_{i}$, $i=1,2$.
+Then we have $X_{1}\leqslant Y_{1}$ a.s.
+under $\mathbb{Q}_{2}$, which further implies $X-X_{1}\geqslant X-Y_{1}$ a.s.
+under $\mathbb{Q}_{2}$.
+By the monotonicity of $\rho$, we have
 $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})$ $\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-Y_{1})$
-$=x_{1}+\rho^{\mathbb{Q}_{2}}((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}})$
+$=x_{1}+\rho^{\mathbb{Q}_{2}}\left((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}\right)$
+$\geqslant\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
+For the case that $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure, let $m_{0}=-\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-X_{1})$.
+Then we have for $x\geqslant\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-X_{1})$
+$\mathbb{Q}_{2}(X-X_{1}\leqslant x)$
+$=\mathbb{Q}_{2}(\{X-X_{1}\leqslant x\}\cap\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\})+\mathbb{Q}_{2}(\{X-X_{1}\leqslant x\}\cap\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\})$
+$\leqslant\mathbb{Q}_{2}(\{X-x_{1}\leqslant x\}\cap\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\})+\mathbb{Q}_{2}(U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1}))$
+$=\mathbb{Q}_{2}(X-Y_{1}\leqslant x),$
+which implies
+$\frac{(F_{X-X_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+}}{\varepsilon}\leqslant\frac{(F_{X-Y_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+}}{\varepsilon},~x\in\mathbb{R}.$
+By the monotonicity of $\rho$ and the fact $\rho$ is an $\varepsilon$-tail risk measure, we have
+$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})$ $\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-Y_{1})$
+$=x_{1}+\rho^{\mathbb{Q}_{2}}\left((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}\right)$
 $\geqslant\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}.$
@@
 Suppose there exists $X_{1}\in\mathcal{X}$ such that $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$.
-Following the same argument as above to show the inverse inequality, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$ and $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$ such that $X_{1}\leqslant Y_{1}$ for all $w\in\Omega$.
+Following the same argument as above to show the inverse inequality, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$ and $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with some $m>m_{0}$ such that $X_{1}\leqslant Y_{1}$ a.s.
+under $\mathbb{Q}_{2}$ if $\mathcal{X}=L^{\infty}$, and $(F_{X-X_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+}\leqslant(F_{X-Y_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+},~x\in\mathbb{R}$ if $\mathcal{X}$ is unbounded and $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure.
 We have
@@
 $=\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
-For $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have $(X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\geqslant y$ a.s.
+If $\mathcal{X}=L^{\infty}$, for $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have $(X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\geqslant y$ a.s.
 under $\mathbb{Q}_{2}$.
 It follows from the monotonicity and law-invariance of $\rho$ that $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant\rho^{\mathbb{Q}_{2}}(y)$.
-Using the monotonicity of $\rho$ again, we have $\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$ Consequently,
+Using the monotonicity of $\rho$ again, we have
+$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
+If $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure, for $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-x)$, we have for $z\geqslant y$
+$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\mathbb{Q}_{2}(\{X-x\leqslant z\}\cap B)+\mathbb{Q}_{2}(B^{c})\leqslant 1=\mathbb{Q}_{2}(y\leqslant z).$
+Moreover, $\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}<y)=\mathbb{Q}_{2}(\{X-x<y\}\cap B)<1-\varepsilon$.
+Hence, $\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant y$ and
+$\frac{(\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)-(1-\varepsilon))_{+}}{\varepsilon}\leqslant\frac{(\mathds{1}_{[y,\infty)]}(z)-(1-\varepsilon))_{+}}{\varepsilon},~z\in\mathbb{R}.$
+Using the monotonicity of $\rho$ and the fact that $\rho$ is an $\varepsilon$-tail risk measure under $\mathbb{Q}_{2}$, we have
+$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
+Consequently,
 $\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}=\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
 Next, we consider the case $\mathbb{P}(B_{1}\cap B_{2}^{c})<(1-\Lambda(x))\mathbb{P}(B_{1})$.
-Let $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$ and $A_{x}=(B_{1}\cap B_{2}^{c})\cup(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})$ with $\alpha_{x}$ satisfying $\mathbb{P}(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})=(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$.
+Let $A_{x}=(B_{1}\cap B_{2}^{c})\cup(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})$ with $\alpha_{x}$ satisfying $\mathbb{P}(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})=(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$.
 For any $B\in\mathcal{F}$ with $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $z\geqslant y$, we have
-$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\frac{\mathbb{P}(X\leqslant z+x,B\cap B_{2})+\mathbb{P}(B^{c}\cap B_{2})}{\mathbb{P}(B_{2})}=1-\frac{\mathbb{P}(X>z+x,B\cap B_{2})}{\mathbb{P}(B_{2})}.$
+$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)$ $=\frac{\mathbb{P}(\{X\leqslant z+x\}\cap B\cap B_{2})+\mathbb{P}(B^{c}\cap B_{2})}{\mathbb{P}(B_{2})}$
+$=1-\frac{\mathbb{P}(\{X>z+x\}\cap B\cap B_{2})}{\mathbb{P}(B_{2})}.$
 Observe that
-$\mathbb{P}(X>z+x,A_{x}\cap B_{2})=\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})\leqslant\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap B_{3})$
+$\mathbb{P}(\{X>z+x\}\cap A_{x}\cap B_{2})=\mathbb{P}(\{X>z+x\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})\leqslant\mathbb{P}(\{X>z+x\}\cap B_{1}\cap B_{2}\cap B_{3})$
 for any $B_{3}\in\mathcal{F}$ such that $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$.
 Note that $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ implies $\mathbb{P}(B_{1}\cap B_{2}\cap B)\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$.
-Hence we have $\mathbb{P}(X>z+x,A_{x}\cap B_{2})\leqslant\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap B)\leqslant\mathbb{P}(X>z+x,B\cap B_{2})$.
+Hence we have $\mathbb{P}(\{X>z+x\}\cap A_{x}\cap B_{2})\leqslant\mathbb{P}(\{X>z+x\}\cap B_{1}\cap B_{2}\cap B)\leqslant\mathbb{P}(\{X>z+x\}\cap B\cap B_{2})$.
 This implies
-$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)\leqslant\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)$
-for $z\geqslant y$. For $z<y$, clearly,
+$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)\leqslant\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z),~z\geqslant y.$
+For $\mathcal{X}=L^{\infty}$, let $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$. Then for $z<y$, clearly,
 $\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)=0.$
-By the law-invariance and monotonicity of $\rho$, we have for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}),$ which implies
+By the law-invariance and monotonicity of $\rho$, for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, and $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have
+$\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}).$
+For the case that $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure, let $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-x)$. Then for $z<y$,
+$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)<1-\varepsilon,$
+which implies $\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant y$. Combing the above conclusion, we have
+$\frac{(\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)-(1-\varepsilon))_{+}}{\varepsilon}\leqslant\frac{(\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)-(1-\varepsilon))_{+}}{\varepsilon},~z\in\mathbb{R}.$
+Using the monotonicity of $\rho$ and the fact that $\rho$ is an $\varepsilon$-tail risk measure, we have for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, and $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-x)$,
+$\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}).$
+Consequently,
 $\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}=\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\}.$
@@
 By Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(x+y)=$-$\infty$.
-Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. Suppose $y<\inf^{\mathbb{Q}_{2}}X-x$. For $z\geqslant x+y$, we have
-$\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)$ $=\frac{\mathbb{P}(A_{x}^{c}\cap B_{2})+\mathbb{P}(X\leqslant z,A_{x}\cap B_{2})}{\mathbb{P}(B_{2})}$ (13)
-$=1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})},$
+Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. Suppose $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X)-x$. For $z\geqslant x+y$, we have
+$\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)$ $=\frac{\mathbb{P}(A_{x}^{c}\cap B_{2})+\mathbb{P}(\{X\leqslant z\}\cap A_{x}\cap B_{2})}{\mathbb{P}(B_{2})}$ (14)
+$=1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})},$
 where $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$.
-Moreover, $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)=0$ for $z<x+y$.
+Moreover, $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)<\varepsilon$ for $z<x+y$.
 If $g(\beta)<1$, then there exists $x\in\mathbb{R}$ such that $g(\beta_{x})<1$.
@@
 Hence, for $x\in\mathbb{R}$ with $\beta_{x}=\beta$, we have
-$\inf_{y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})=\lim_{y\to-\infty}\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)=-\infty.$
-Now, we assume $g(\beta)=1$. Direct computation shows for $0<t<\beta_{x}$,
-$\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})$ $=\inf\{z\geqslant x+y:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}\geqslant 1-t\}$
+$\inf_{y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\leqslant\lim_{y\to-\infty}\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)=-\infty.$
+Now, we assume $g(\beta)=1$. Direct computation shows for $0<t<\beta_{x}\wedge\varepsilon$,
+$\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})$
+$=\inf\left\{z\geqslant x+y:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}\geqslant 1-t\right\}$
 $=\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X).$ (15)
 Hence, we have
-$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\int_{[0,\beta)}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)$
-$=\inf_{x\in\mathbb{R}}\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t)=\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
-Let us now consider (ii).
-If $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then there exists $x\in\mathbb{R}$ such that $\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})$.
-Hence, $A_{x}=\emptyset$.
-By Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+u(y))=-\infty$.
-Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. It follows from Proposition 5 and (C) that
-$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{2}}u((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\}$
-$=\inf_{x\in\mathbb{R}}\left(x+(1-\beta_{x})u(-\infty)+\int_{0}^{\beta_{x}}u\left(\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)-x\right)\mathrm{d}t\right),$
-where $u(-\infty)=\lim_{y\to-\infty}u(y)$.
-Finally, we consider (iii).
+$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\int_{[0,\beta_{x}\wedge\varepsilon)}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)$
+$=\inf_{x\in\mathbb{R}}\int_{[0,\beta_{x}\wedge\varepsilon)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t)=\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
+Let us now consider case (ii).
 Similarly as above, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then by Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+y)=$-$\infty$.
@@
 $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)$
-$=\inf_{x\in\mathbb{R}}\lim_{y\to-\infty}\left\{x+\sup\{z\geqslant y:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant x+z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z)\}\right\}$
-$=\inf_{x\in\mathbb{R}}\inf_{y\to-\infty}\sup\left\{z\geqslant y+x:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
-$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
-$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<\beta_{x}-\Lambda_{1}(z-x)\right\}$
-$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{1}\cap B_{2})}<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}$
+$=\inf_{x\in\mathbb{R}}\lim_{y\to-\infty}\left\{x+\sup\{z\geqslant y:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant x+z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z)\}\right\}$
+$=\inf_{x\in\mathbb{R}}\inf_{y\to-\infty}\sup\left\{z\geqslant y+x:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
+$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
+$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<\beta_{x}-\Lambda_{1}(z-x)\right\}$
+$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2})}{\mathbb{P}(B_{1}\cap B_{2})}<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}$
 $=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:F_{X}^{\mathbb{Q}}(z)<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}=\inf_{x\in\mathbb{R}}\overline{\Lambda}^{x}\mathrm{VaR}^{+,\mathbb{Q}}(X),$
@@
 $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\inf_{t\in\mathbb{R}}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}((X-t)_{+}\mathds{1}_{B}+(x+y-t)_{+}\mathds{1}_{B^{c}})\right\}$
+$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\inf_{t\in\mathbb{R}}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}((X-t)_{+}\mathds{1}_{B})\right\}$
+$=\inf_{x\in\mathbb{R}}\inf_{t\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{1}}(\eta(X-t)_{+}\mathds{1}_{B})+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}((X-t)_{+}\mathds{1}_{B\cap N})\right\}$
 $=\inf_{x\in\mathbb{R}}\inf_{t\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{1}}(\eta(X-t)_{+}\mathds{1}_{B})\right\}$
@@
 $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$
-$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{B})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B^{c}})\right\}$
-$=\inf_{x\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})+u(-\infty)\int_{0}^{\Lambda(x)}\mathrm{VaR}_{t}^{\mathbb{Q}_{1}}(\eta)\mathrm{d}t\right\}.$
-Appendix D Proof of Section 5
-In this section, we offer all the proofs of the results in Section 5.
+$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{B})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B^{c}})\right$.
+$~~~~\left.+\mathbb{E}^{\mathbb{Q}_{2}}(u(X-x)\mathds{1}_{B\cap N})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{2}}(\mathds{1}_{B^{c}\cap N})\right\}$
+$=\inf_{x\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})+u(-\infty)\left(\mathbb{Q}_{2}(N)+\int_{0}^{\Lambda(x)}\mathrm{VaR}_{t}^{\mathbb{Q}_{1}}(\eta)\mathrm{d}t\right)\right\}.$
 Proof of Theorem 6. For $x\in\mathbb{R},y\geqslant x$ and $U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]$, we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(\Lambda_{x,y}^{-1}(U))=x$. Hence,
@@
 This implies $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for $x_{1}\leqslant x\leqslant y$.
+Suppose $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}<\infty$.
 If $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, then $F_{X_{1}}^{\mathbb{Q}_{1}}(x)=1\geqslant\Lambda_{x_{1},y}(x)$ for $x\geqslant y$.
@@
 $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
+If $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}=\infty$, then let $X_{m}=X_{1}\wedge m$ for $m\geqslant 2$. Using the above conclusion, we have
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{m})+\rho^{\mathbb{Q}_{2}}(X-X_{m})\geqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
+Direct computation gives $\lim_{m\to\infty}\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{m})=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$.
+Note that $X-X_{m}\downarrow X-X_{1}$ as $m\to\infty$.
+Using the fact that $\rho$ is continuous from above, we have $\lim_{m\to\infty}\rho^{\mathbb{Q}_{2}}(X-X_{m})=\rho^{\mathbb{Q}_{2}}(X-X_{1})$.
+Hence, letting $m\to\infty$, we obtain
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})\geqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
 By the arbitrary of $X_{1}$, we have
@@
 Hence, we obtain (10).
-Next, we show the optimal allocation of the inf-convolution exists if and only if the minimizer of (9) exists.
+Next, we show for $\mathcal{X}=L^{\infty}$, the optimal allocation of the inf-convolution exists if and only if the minimizer of (10) exists.
 Suppose $(X_{1},X-X_{1})$ is the optimal allocation of the inf-convolution.
@@
 ∎
+Proof of Corollary 8. In light of Theorem 6, we have
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}\mathop{\square}\displaylimits\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\left\{x+\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
+Let $\hat{F}_{x,y}(z)=\{F^{\mathbb{Q}}_{X-\Lambda_{x,y}^{-1}(U)}(z):U\overset{\mathbb{Q}}{\sim}U[0,1]\}$ for $z\in\mathbb{R}$ and $\overline{\mathrm{VaR}}_{\beta}^{\mathbb{Q}}$ represent $\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}$ for $\Lambda_{1}=\beta\in(0,1)$.
+This means $\overline{\mathrm{VaR}}_{\beta}^{\mathbb{Q}}$ is the right $\mathrm{VaR}$.
+By Theorem 1 of Han and Liu (2025), we have
+$\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U)\right)=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y}).$
+It follows from Corollary 1 of Han and Liu (2025) that for $\alpha\in(0,1)$,
+$q_{\alpha}^{+}(\hat{F}_{x,y})=\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{\alpha}^{\mathbb{Q}}(X-\Lambda_{x,y}^{-1}(U)),$
+where $q_{\alpha}^{+}$ is $q_{\Lambda_{1}}^{+}$ with $\Lambda_{1}=\alpha$. Direct computation shows
+$\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{\alpha}^{\mathbb{Q}}(X-\Lambda_{x,y}^{-1}(U))=-\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\mathrm{VaR}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U)).$
+In light of Lemma 4.3 of Bernard et al. (2014) and the result in Makarov (1981), we have
+$\sup_{Y\overset{\mathbb{Q}}{\sim}F_{X}^{\mathbb{Q}},U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-Y+\Lambda_{x,y}^{-1}(U))$ $=\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(V))$
+$\leqslant\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U)),$
+where $V=U_{-X}^{\mathbb{Q}}\mathds{1}_{\{U_{-X}^{\mathbb{Q}}\leqslant\alpha\}}+(1+\alpha-U_{-X}^{\mathbb{Q}})\mathds{1}_{\{U_{-X}^{\mathbb{Q}}>\alpha\}}$, which implies
+$\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U))$ $=\sup_{Y\overset{\mathbb{Q}}{\sim}F_{X}^{\mathbb{Q}},U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-Y+\Lambda_{x,y}^{-1}(U))$
+$=\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(V)).$
+Using the above conclusion and the continuity of $\Lambda_{x,y}^{-1}$ and $F_{X}^{\mathbb{Q},-1}$, applying Lemma 4.4 of Bernard et al. (2014), we have
+$\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\mathrm{VaR}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U))=\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U)).$
+Hence,
+$q_{\alpha}^{+}(\hat{F}_{x,y})=-\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(V))=l_{x,y}(1-\alpha),$
+where $l_{x,y}(t)=\sup_{s\in(0,t]}\{F_{X}^{\mathbb{Q},-1}(s)-\Lambda_{x,y}^{-1}(1-t+s)\},~t\in(0,1)$.
+Direct computation shows $\hat{F}_{x,y}(z)=l_{x,y}^{-1}(z)$ on $\mathbb{R}$ except a countable number of points.
+Next, we show $q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$.
+Let $z_{0}=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$.
+Clearly, $z_{0}\in\mathbb{R}$.
+By definition, there exist $z_{n}\downarrow z_{0}$ such that $\hat{F}_{x,y}(z_{n})>1-\Lambda_{1}(z_{n})$.
+It follows from the right continuity of $\Lambda_{1}$ and the monotonicity of $\hat{F}_{x,y}$ that there exist $0<\varepsilon_{n}<1/n$ such that $\hat{F}_{x,y}(z)>1-\Lambda_{1}(z)$ for $z\in(z_{n},z_{n}+\varepsilon_{n})$.
+Then there exist $w_{n}\in(z_{n},z_{n}+\varepsilon_{n})$ such that $l_{x,y}^{-1}(w_{n})=\hat{F}_{x,y}(w_{n})>1-\Lambda_{1}(w_{n})$.
+Hence, $q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})\leqslant w_{n}\to z_{0}=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$.
+We can similarly show the inverse inequality.
+Hence, $q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$.
+We complete the proof.
+∎
 Proof of Proposition 6.
@@
 $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-x\mathds{1}_{B_{1}}-x\vee x_{1}\mathds{1}_{B_{2}}-y\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)\right\}.$
+Proof of Corollary 9. In light of Proposition 6, we have
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\left\{x+\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-x\mathds{1}_{B_{1}}-x\vee x_{1}\mathds{1}_{B_{2}}-y\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)\right\}$
+$=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X+0\wedge(x-x_{1})\mathds{1}_{B_{2}}+(x-y)\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)$
+$=\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\lim_{y\to\infty}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-y\mathds{1}_{B_{1}^{c}}\right).$
+From the above expression, we should choose $B_{1}$ such that $\mathbb{Q}_{2}(B_{1})$ is as small as possible.
+Hence, if $\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})\geqslant\alpha\mathbb{P}(D_{2})$, the optimal way to choose $B_{1}$ is to set $B_{1}=(D_{1}\cap D_{2}^{c})\cup(\{U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})$ with $\mathbb{P}(\{U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})=\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})$.
+Then we have
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\lim_{y\to\infty}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-y\mathds{1}_{B_{1}^{c}}\right)=\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}\left(X\mathds{1}_{\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2}}\right)$
+with $\mathbb{P}(\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})=\alpha\mathbb{P}(D_{2})$.
+Note that $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}}=\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}$ on $D_{1}\cap D_{2}$.
+Hence, we have
+$\mathbb{E}^{\mathbb{Q}_{2}}\left(X\mathds{1}_{\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2}}\right)$ $=\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}\mathbb{E}^{\mathbb{Q}}\left(X\mathds{1}_{\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2}}\right)$
+$=\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}\int_{\alpha_{0}}^{\alpha_{1}}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t.$
+with $\alpha_{0}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})}{\mathbb{P}(D_{1}\cap D_{2})}$ and $\alpha_{1}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})-\alpha\mathbb{P}(D_{2})}{\mathbb{P}(D_{1}\cap D_{2})}$, which implies the desired result.
+If $\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})\leqslant 0$, we choose $B_{1}\subset D_{1}\cap D_{2}^{c}$.
+Then $\mathbb{Q}_{2}(B_{1}^{c})=1$, which implies $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=$-$\infty$.
+If $0<\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})<\alpha\mathbb{P}(D_{2})$, then we choose $B_{1}=(D_{1}\cap D_{2}^{c})\cup(\{U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})$, where $t^{+}$ is defined above.
+Direct computation shows $\mathbb{Q}_{2}(B_{1})<\alpha$, which implies
+$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\lim_{y\to\infty}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-y\mathds{1}_{B_{1}^{c}}\right)=-\infty.$
+This completes the proof. ∎